### v2

# The Type Ia Supernova 2021hem: A 2003fg-like Event in an Apparently Hostless Environment

11institutetext: Department of Physics and Astronomy, Aarhus University, Ny Munkegade 120, DK-8000 Aarhus C, Denmark (11email: email@subhashbose.com, max@phys.au.dk) 22institutetext: Department of Astronomy, University of Michigan, Ann Arbor, MI 48109, USA 33institutetext: INAF – Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy 44institutetext: Cosmic DAWN Center, Niels Bohr Institute, University of Copenhagen, Lyngbyvej 2, 2100 Copenhagen Ø, Denmark 55institutetext: Institute for Astronomy, University of Hawai’i, 2680 Woodlawn Drive, Honolulu HI 96822, USA 66institutetext: Observatories of the Carnegie Institution for Science, 813 Santa Barbara St, Pasadena, CA, 91101, USA 77institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218-2410, USA 88institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, E-08193 Barcelona, Spain 99institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), E-08034 Barcelona, Spain 1010institutetext: Department of Physics, Florida State University, 77 Chieftan Way, Tallahassee, FL 32306, USA 1111institutetext: Department of Physics, Chalmers University of Technology, SE-412 96, Göteborg, Sweden 1212institutetext: European Southern Observatory, Alonso de Córdova 3107, Casilla 19, Santiago, Chile 1313institutetext: Planetary Science Institute, 1700 East Fort Lowell Road, Suite 106, Tucson, AZ 85719-2395, USA 1414institutetext: CENTRA, Instituto Superior Técnico, Universidade de Lisboa, Av. Rovisco Pais 1, 1049-001 Lisboa, Portugal 1515institutetext: Astronomical Observatory, University of Warsaw, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 1616institutetext: Cardiff Hub for Astrophysics Research and Technology, School of Physics & Astronomy, Cardiff University, Queens Buildings, The Parade, Cardiff, CF24 3AA, UK 1717institutetext: Astrophysics Research Institute, Liverpool John Moores University, IC2, Liverpool Science Park, 146 Brownlow Hill, Liverpool L3 5RF, UK 1818institutetext: School of Physics, Trinity College Dublin, The University of Dublin, Dublin 2, Ireland 1919institutetext: Instituto de Ciencias Exactas y Naturales (ICEN), Universidad Arturo Prat, Chile 2020institutetext: The Oskar Klein Centre, Department of Physics, Stockholm University, AlbaNova, 10691 Stockholm, Sweden 2121institutetext: INAF - Osservatorio Astronomico di Capodimonte, Salita Moiariello 16, 80131 Napoli, Italy 2222institutetext: Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721-0065, USA
The Type Ia Supernova 2021hem: A 2003fg-like Event in an Apparently Hostless Environment
Subhash Bose[Uncaptioned image]    M. D. Stritzinger[Uncaptioned image]    A. Malmgaard    C. J. Miller[Uncaptioned image]    N. Elias-Rosa[Uncaptioned image]    J. P. U. Fynbo[Uncaptioned image]    C. Ashall[Uncaptioned image]    C. R. Burns[Uncaptioned image]    J. M. DerKacy[Uncaptioned image]    L. Galbany[Uncaptioned image]    C. P. Gutiérrez[Uncaptioned image]    W. B. Hoogendam[Uncaptioned image]    E. Y. Hsiao[Uncaptioned image]    E. A. M. Jensen[Uncaptioned image]    K. Medler[Uncaptioned image]    Alaa Alburai[Uncaptioned image]    J. Anderson[Uncaptioned image]    E. Baron[Uncaptioned image]    J. Duarte[Uncaptioned image]    M. Gromadzki[Uncaptioned image]    C. Inserra[Uncaptioned image]    P. A. Mazzali[Uncaptioned image]    T. E. Müller-Bravo[Uncaptioned image]    P. Lundqvist    A. Reguitti[Uncaptioned image]    I. Salmaso[Uncaptioned image]    D. J. Sand[Uncaptioned image]    G. Valerin[Uncaptioned image]
(Received 10 November 2025 / Accepted 24 November 2025.)
We report observations of a Type Ia supernova 2021hem, discovered within 48 hours of last non-detection, and located in an apparently hostless environment. With a peak absolute $B$-band magnitude of $M_{B,\max}=-19.96\pm 0.29$ mag, SN 2021hem lies at the luminous end of the SN Ia distribution. Its near-infrared and $i$-band light curves lack the secondary maximum, which is otherwise ubiquitous to normal and 1991T-like SNe Ia. Instead, these properties make SN 2021hem closely resemble 2003fg-like events. The slowly evolving light curves (characterized by $\Delta m_{15}(B)=1.02\pm 0.02$ mag; $s_{BV}=0.94\pm 0.05$), and the earliest spectrum showing C ii $\lambda$6580 and $\lambda$7235 absorption lines, further support this classification. Other spectroscopic features, including Si ii line diagnostics, resemble normal SNe Ia. A fit of a “fireball” model to the early-time light curves yields a time of first light of $t_{\rm first}=-16.43^{+0.45}_{-0.38}$ days relative to $B$-band maximum. The first photometric detection occurs $1.51^{+0.45}_{-0.38}$ days before the onset of fireball-like flux rise. This early emission, together with the intrinsic $(g-r)_{0}$ color, is inconsistent with circumstellar or companion interaction. Instead, shallow 56Ni mixing or an asymmetric 56Ni distribution offers a plausible explanation for the delayed onset of the fireball flux rise, while a double-detonation scenario with a thin helium shell remains a less likely alternative. Notably, SN 2021hem represents the fifth known 2003fg-like SN that has early-time activity or excess flux emission. The estimated mass of radioactive 56Ni synthesized in SN 2021hem is $1.00\pm 0.09$ M. Deep GTC imaging obtained 2.5 years after the explosion, with an estimated limiting magnitude of $m_{lim,r}=24.4$ mag and a surface-brightness limit of $\mu_{lim,r}=26.3$$\mathrm{mag\penalty 10000\ arcsec^{-2}}$, reveals no coincident host, thereby ruling out most faint dwarf and ultra-diffuse galaxies (UDGs). Alternatively, if the nearest plausible AGN host galaxy located at a projected distance of 104  kpc is assumed, the progenitor would need to be a hyper-velocity star ejected at $\sim 2200$$\rm{\,km\,s^{-1}}$ from the host by AGN interaction. A faint diffuse feature $\approx 6$ kpc from the SN site has also been detected in the GTC image, with its surface brightness within the limits of UDGs. However, it is unclear whether it is a galaxy and is associated with SN 2021hem. Considering its large normalized directional light distance ($d_{\mathrm{DLR}}$ $\sim 3-4$) from the SN, and its unusual elongation, it is a candidate of low probability to be the host galaxy of SN 2021hem. These results identify SN 2021hem as one of the strongest candidates for a hostless SN Ia, underscoring the diversity of luminous, slowly evolving, 2003fg-like explosions and the wide range of environments in which they may occur.
Key Words.:
supernovae : general – supernovae: individual: SN 2021hem
1 Introduction
Type Ia supernovae (SNe Ia) are thermonuclear explosions of carbon–oxygen white dwarfs (WDs), generally thought to occur when the progenitor approaches the Chandrasekhar mass limit of $\approx\penalty 10000\ 1.4$$M_{\odot}$ (Nomoto1984; Hillebrandt2000). Their relative uniformity in peak luminosity has made them indispensable as cosmological distance indicators (Riess1998; Perlmutter1999). Beyond the population of normal SNe Ia, several distinct subclasses have been identified, including the low-luminosity, red 1991bg-like events (e.g., 1992AJ....104.1543F; 1993AJ....105..301L), the overluminous, blue 1991T-like SNe (e.g., 1992ApJ...384L..15F; 1992AJ....103.1632P; 2024ApJS..273...16P), and the exceptionally bright 2003fg-like explosions (Howell2006). Additional peculiar subtypes have also been recognized, such as the 2002cx-like (commonly referred to as SNe Iax; 2003PASP..115..453L; 2013ApJ...767...57F) and the 2002es-like SNe (e.g., 2012ApJ...751..142G; Bose2025).
Among the rarest subtypes of SNe Ia are the 2003fg-like events (Howell2006; 2007ApJ...669L..17H), historically referred to as “super-Chandrasekhar” SNe Ia. These objects exhibit broad, slowly declining light curves and peak luminosities exceeding those of normal SNe Ia with comparable decline rates. They also lack the secondary maximum in near-infrared (NIR) light curves that typically characterizes normal or luminous 1991T-like SNe, while often displaying several distinctive features. These include: low expansion velocities, a prominent C ii $\lambda$6580 absorption feature near maximum light, and early light curves that show a short-lived flux excess, possibly arising from interaction with circumstellar material. Furthermore, their relatively low expansion velocities, despite their high luminosities, are consistent with large ejecta masses (Howell2006; 2011MNRAS.412.2735T; Scalzo2010). This led to the interpretation that they originate from super-Chandrasekhar-mass WD progenitors. However, some 2003fg-like SNe have inferred ejecta masses within the Chandrasekhar-mass limit (see 2014MNRAS.443.1663C; Lu2021), suggesting a diversity in progenitor systems. The requirement for high ejecta masses in these events could also be satisfied by the merger of two near-Chandrasekhar-mass WDs (e.g., 2023MNRAS.521.1162D). Alternatively, CSM-interaction models have been invoked to explain the high luminosity in several 2003fg-like SNe (e.g., 2012MNRAS.427.2057H; 2020ApJ...900..140H; 2024A&A...687L..19N; 2025MNRAS.542.2752B). Many 2003fg-like SNe also show a tendency to occur in low-mass, often metal-poor, dwarf host galaxies (Childress2011; Khan2011). Although studies based on larger samples indicate that such events can arise in hosts spanning a wide range of masses (2025A&A...694A..10D; 2025MNRAS.542.2752B).
SNe Ia occur in a wide range of galactic environments, extending from star-forming spirals to quiescent ellipticals, and even within dense galaxy clusters, where intracluster stars might serve as progenitors (2006ApJ...648..868S; Lampeitl2010; Pan2014). While the majority of SNe Ia are clearly associated with galaxies, a small fraction appear to lack a host-galaxy association. Only a handful of confirmed SNe Ia to date have been identified in the intracluster regions of galaxy clusters, located far from any cluster members and likely originating from the intracluster stellar population (2003AJ....125.1087G; 2011ApJ...729..142S; 2015ApJ...807...83G). In principle, “hostless” SNe could also arise from hyper-velocity stars that escaped their host galaxies prior to explosion. However, no confirmed examples of this scenario have been reported to date. Many candidates labeled as hostless in imaging surveys may, in fact, be associated with low surface brightness dwarf galaxies that remain undetected at the survey’s imaging depth (e.g., Strolger2002; 2025ApJ...988..278S). This underscores the need for deep imaging to robustly confirm truly hostless SNe.
The Type Ia SN 2021hem (ZTF21aaqwjlz) provides an important case of this phenomenon. The absence of a clear host galaxy or any nearby cluster in its field raises the question of whether the explosion was truly hostless or associated with a system below current detection limits. This possibility connects naturally to the population of low-surface-brightness galaxies. Ultra-diffuse galaxies (UDGs), with effective radii of $r_{\rm eff}\gtrsim 1.5$ kpc, having low stellar masses of $\sim 10^{8}\,\mbox{M${}_{\odot}$}$, and central surface brightnesses of $\sim 24$–26.5 mag arcsec-2, have been identified in clusters, groups, and lower-density environments (vanDokkum2015; Koda2015; Munoz2015; Mihos2015; vanderBurg2016; Ramon2017; Janssens2017; 2017A&A...607A..79V). At even fainter extremes lie the ultra-faint dwarf galaxies (UFDGs), with luminosities as low as $M_{V}\sim-1$ to 8-8 and surface brightnesses $\mu_{V}\sim 27$–31 mag arcsec-2, making them among the most diffuse stellar systems known (Willman2005; Belokurov2007; McConnachie2012; DrlicaWagner2015; Torrealba2019; 2019ARA&A..57..375S; Collins2022). They are typically comparable in size to globular clusters, with radii ranging from 30 to 70 pc, although a few larger ones can extend to a few hundred parsecs. Unlike globular clusters, however, UFDs contain several orders of magnitude fewer stars and exhibit higher velocity dispersions of 2–10 $\rm{\,km\,s^{-1}}$, making them among the most dark-matter dominated galactic systems (2023A&A...679A...2R). Due to their low surface brightness and small sizes, UFDGs cannot be detected beyond $\sim$10 Mpc, even with 10-m class telescopes under arcsecond seeing conditions. The most distant UFDG identified so far, found in HST imaging, lies at $\sim$19 Mpc (2017ApJ...835L..27L). Nevertheless, owing to their extreme faintness, these galaxies would always remain plausible hosts for otherwise “hostless” supernovae.
In this paper, we report follow-up data on SN 2021hem, obtained from shortly after explosion through over a month beyond peak brightness, in an apparently hostless environment. This case illustrates the challenges of identifying the faintest hosts and probing the extremely low-luminosity galaxy population. Recent results from the ELEPHANT project indicate that fewer than $\sim$2% of extragalactic transients are potentially hostless (Pessi2024). This highlights the need for deeper surveys to identify such rare events with better certainty.
The paper is organized as follows. Sect. 2 summarizes discovery and classification. Sect. 3 describes the photometric and spectroscopic data sets, reductions, deep late-time imaging, and reddening and redshift estimates. Sect. 4 presents the light-curve analysis, spectroscopic diagnostics. Sect. 5 details the search for an underlying host, deriving point-source and surface-brightness limits and noting a nearby diffuse feature. Finally, Sect. 6 concludes with a brief summary.
2 Discovery and classification
SN 2021hem was identified by ALeRCE (Automatic Learning for the Rapid Classification of Events) broker system with J2000 coordinates R.A. $16^{h}21^{m}16^{s}.01$ and Decl. $14^{\circ}33^{\prime}09\aas@@fstack{\prime\prime}61$ (Forster2021). This discovery was made possible with ALeRCE’s stamp classifier (Carrasco-Davis2021), utilizing data from the Zwicky Transient Facility (ZTF; Bellm2019) public data stream. According to the Transient Name Server (TNS) discovery certificate, the first detection of the object appeared in an $r_{ZTF}$-band image taken on 2021-03-24T10:05:20.00 UT and the last non-detection was obtained 45 hours earlier (i.e, 2021-03-22T11:33:32 UT). Making use of the ZTF forced photometry service (Masci2023), we recomputed photometry of the ZTF data stream which indicates at the time of discovery of SN 2021hem had an $r_{ztf}$-band magnitude of 19.9, while the limiting magnitude of the non-detection image was 20.12 mag.
An optical spectrum obtained with the 6.5-m MMT Observatory equipped with Binospec on 2021-04-01T09:33:20 allowed Terreran2021 to classify SN 2021hem as a SN Ia. They also reported a redshift of $z=0.035$, presumably inferred from spectral template comparisons.
Figure 1: $r$-band image of SN 2021hem at peak brightness observed with the Nordic Optical Telescope. The location of SN 2021hem is positioned at the center in both panels. Two of the closest cataloged galaxies are marked in blue, both at the projected distance of roughly $\sim 100\mbox{\penalty 10000\ kpc}$ (see Sect. 5.1)
Based on the initial ZTF public stream data and demonstrated by the finding chart provided in Fig. 1, SN 2021hem appears to lack an associated host galaxy. This motivated our team to conduct a followup campaign covering the early photospheric phase evolution.
Figure 2: Large-scale structure around the position of SN 2021hem (marked with a cyan cross) shown in a field of a size of $\times2$ square degrees DSS image. The colored circles indicate the galaxies with measured redshifts from SIMBAD catalog within redshift range of 0.02 to 0.05. The color-bar indicates the redshifts for each color. There appears to be an overdensity of galaxies with redshifts close to $z\sim 0.027$, $\sim 0.035$ within a radius of 1 Mpc from the position of the SN. The open star symbol corresponds to the AGN galaxy 2MASX J16212572+1431537, which has the projected distance of 104 kpc from the SN.
3 Observations
3.1 SN data
Our characterization of the photometric properties of SN 2021hem are based on a collection of photospheric phase optical $BgVri$- and near-IR $JH$-band photometry. During the course of the followup campaign, 22 epochs of optical imaging were obtained with the Las Cumbres Observatory’s Global Network of 1-m robotic telescopes and a single epoch with the Nordic Optical Telescope (NOT) equipped with the Alhambra Faint Object Spectrograph and Camera (ALFOSC). Two epochs of the $JH$-band imaging were also obtained with the NOT, equipped with the NOT near-infrared Camera and spectrograph (+ NOTcam), while an additional epoch of $JH$-band imaging was obtained with the New Technology Telescope (NTT) equipped with the Son OF ISAAC (Sofi).
The optical and NIR imaging data sets were reduced following standard procedures. PSF photometry was computed for the SN using the FLOWS [^1] https://flows.phys.au.dk photometry pipeline, with nightly zeropoints computed relative to stars in the field with ATLAS Refcat2 catalog magnitudes (Tonry2018b).
Our Las Cumbres photometry is complemented by the ZTF public data stream, and Pan-STARRS photometry through Young Supernova Experiment (YSE; 2021ApJ...908..143J). The ZTF forced photometry service was used to compute up to date photometry and we ensured the implementation of the baseline correction (see Masci2023, for details). This data provides high cadence photometry beginning from around the time of first light (hereafter $t_{first}$; see Sect. 4.1.2) and extending though the primary maximum. Finally, for completeness, we also ingested $o$- and $c$-band forced photometry from The Atlas Project (Tonry2018a; Shingles2021). The photometry measurements are listed in Tables LABEL:tab:photsn-LABEL:tab:photsnNIR.
Table 1: Journal of Spectroscopic Observations.222
Date Julian Date PhaseaaaaaaRestframe days relative to the epoch of $B$-band maximum, i.e., JD 2459316.46. Telescope Instrument Integration (s) $z$ rlapbbbbbbEach redshift inferred by SNID carries a rlap parameter. This is a reliability metric that reflects both the strength of the spectral cross-correlation and the extent of wavelength overlap between the observed spectrum and the matched template. Higher rlap values correspond to better spectral matches, with values above $\sim 5$ generally considered significant for classification.
Optical
2021 April 01ccccccTerreran2021 reported a redshift of $z=0.035$ to the Transient Name Server along with their classification spectrum. 2459305.90 -10.19 MMT BINOSPEC 300 $0.030\pm 0.006$ 9.6
2021 April 10 2459314.57 -1.82 NOT ALFOSC 900 $0.037\pm 0.005$ 12.1
2021 April 13 2459317.90 ++1.39 APO 3.5-m DIS 1800 $0.036\pm 0.006$ 14.5
2021 April 16 2459320.56 ++3.96 NOT ALFOSC 900 $0.035\pm 0.003$ 18.1
2021 April 20 2459324.62 ++7.87 NOT ALFOSC 900 $0.044\pm 0.004$ 16.0
2021 May 05 2459338.69 ++21.45 NOT ALFOSC 900 $0.035\pm 0.004$ 15.0
2021 May 16 2459351.50 ++33.81 NOT ALFOSC 1800 $0.035\pm 0.003$ 17.3
NIR
2021 April 11 2459315.91 0.53-0.53 IRTF SpeX 1079 $\cdots$ $\cdots$
In addition to the classification spectrum available on TNS (Transient Name Server), we obtained, in coordination with NUTS2 (NOT Unbiased Transient Survey), 5 optical spectra with the NOT (+ ALFOSC), as well as a single spectrum with the APO 3.5-m (+ DIS, dual imaging spectrograph). The NOT data were reduced following standard procedures using a software graphical user interface written in pyraf. Nightly sensitivity functions were computed based on the observations of at least one nightly spectrophometric flux standard. Similar reductions were performed using IRAF packages on the APO data obtained with both red and blue gratings. Complementing the optical spectroscopic time-series is a single NIR spectrum of SN 2021hem obtained with the NASA IRTF (+ SpecX) as part of The Hawaii Infrared Supernova Study (2025arXiv250518507M) at around peak brightness. A journal of spectroscopic observations is provided in Table 1.
3.2 Late-time deep imaging by GTC
To characterize the location of SN 2021hem, we obtained a sequence of $r$-band images with the Gran Telescopio CANARIAS (GTC) equipped with OSIRIS (Optical System for Imaging and low-Intermediate-Resolution Integrated Spectroscopy), $\sim 900$ days after the explosion on 9 September 2023. At this phase, the SN is expected to be around $\sim 27-28$ mag assuming SN 2011fe like light curve, which is well below the detection limit of the image (see Sect. 5.2.1 for the precise estimate of the image limiting magnitude). In total, 7 individual images of 60 second integration time were obtained, along with a single image with a 15 second integration time.
The images were reduced using standard procedures with software based on the photutils package. Each image was bias-subtracted, flat-field corrected, and cleaned of cosmic rays. The images were then geometrically aligned and combined to produce a single science frame. This image is used in Sect. 5.2.1 to estimate the limiting magnitude of any potential underlying host galaxy.
3.3 Reddening
The Milky Way reddening color excess along the line-of-sight toward SN 2021hem is $E(B-V)_{MW}=0.0501\pm 0.0034$ mag (2011ApJ...737..103S). The post-peak $(B-V)_{0}$ color curve of SN 2021hem, after correcting for Milky way reddening is consistent with the intrinsic color-curve of the archetypal normal SN 2011fe as well as the subluminous SN 1991bg (see Fig. 12), suggesting SN 2021hem suffered minimal host extinction. Further corroborating this, SN 2021hem spectra reveal no Na i D absorption at the redshift of the SN. The minimal host reddening is also consistent with the apparent hostless environment of the SN. Therefore, we adopt the previously stated Milky Way reddening value in our extinction correction for SN 2021hem. Assuming $R_{V}=3.1$, this corresponds to an $A^{tot}_{V}=0.155\pm 0.011$ mag.
3.4 Redshift of SN 2021hem
Using the Supernova Identification (SNID) code (Tonry1979; Blondin2007), we determined the redshift of SN 2021hem for each spectrum by comparing it to a library of supernova spectral templates. The individual redshift estimates, their associated uncertainties, and the corresponding rlap values – which quantify the quality of the spectral matches – are listed in Table 1. We computed a weighted mean redshift incorporating both the statistical uncertainties and the rlap values as quality weights, yielding a redshift of $z=0.0363\pm 0.0044$. To account for potential systematic uncertainties in the SNID-based redshift estimates, arising from template mismatch, phase coverage limitations, and spectral quality, a systematic error floor of $\Delta z_{sys}=0.002$ is adopted (Blondin2007; Foley2009), which was added in quadrature to the statistical uncertainty. This yields a redshift estimate of $z=0.0363\pm 0.0049$. This value is adopted in the following to apply redshift corrections and to compute the corresponding luminosity distance of $D_{L}=165\pm 22$ Mpc assuming a standard Planck cosmology (2016A&A...594A..13P), and the equivalent distance modulus of $36.09\pm 0.29$ mag. We note the potential bias in SNID template matching for a low velocity SN, which we discuss in detail in Sect. 4.3, and conclude that SN 2021hem is a normal velocity SN Ia and unlikely there is any significant bias in the determined redshift.
As a consistency check, we examine the redshift distribution of galaxies within the field of view of SN 2021hem, as shown in Fig. 2. The distribution reveals over-densities at redshifts of $z\sim 0.027$, $\sim 0.035$, as well as a weaker concentration at $\sim 0.050$, suggesting the presence of three major galaxy groups along the line-of-sight to SN 2021hem. Given the redshift determined from SNID template matching, $z=0.0363$, SN 2021hem is plausibly associated with the galaxy group at $z\sim 0.035$.
Figure 3: Optical and NIR light curves of SN 2021hem, shown alongside the best-fit SNooPy template light curves using the ‘max_model’. Our own observations are shown with filled circles, while photometry from ZTF, YSE, and ATLAS surveys are shown in different symbols. The SNooPy template fits are overplotted as solid lines; those in the grigri bands are derived specifically from fits to the ZTF photometry. The indicated offsets have been added to the light curves for clarity. The vertical dashed line corresponds to the estimated “time of first light” by fitting a “fireball” model (see Sect. 4.1.2).
4 Results
4.1 Photometry
4.1.1 Light-curve properties and template fitting
The optical and NIR light curves of SN 2021hem, plotted relative to the epoch of $B$-band maximum, are shown in Fig. 3. The $B$-band maximum is estimated to have occurred on $\rm JD\penalty 10000\ 2459316.46\pm 0.31$, based on a direct Gaussian Process spline fit to the $B$-band light curve. This has been used as the reference epoch throughout the paper.
Using spline fits, the estimated peak apparent magnitude is $m_{B,max}=16.34\pm 0.01$ mag, and the post-peak light curve decline is estimated to be $\mbox{$\Delta m_{15}(B)$}=1.02\pm 0.02$ mag. After correcting for Milky Way reddening, computed K-correction of 0.005-0.005 mag, and luminosity distance for the adopted redshift of $z=0.0363\pm 0.0049$, the peak absolute magnitude of SN 2021hem is $M_{B,max}=-19.96\pm 0.29$. On constructing $(B-V)_{0}$ intrinsic color-curve with interpolated spline fits after correcting for reddening, the color-curve peaks at $28.3\pm 1.1$ day, which corresponds to color-stretch parameter (2014ApJ...789...32B) $\mbox{$s_{BV}$}=0.94\pm 0.05$.
In Fig. 3, the observed light curve data points are overlaid with template fits generated using the SNooPy package (Burns2011), adopting the ‘max_model’ formalism. The ‘max_model’ in SNooPy fits the peak magnitude in each band directly by treating the maximum light in each filter as free parameters, while using the same template-driven time evolution (based on $\Delta m_{15}$) and applying $K$-corrections and Milky Way extinction consistent with the ‘EBV_model’ framework. Unlike models that impose an explicit reddening law and intrinsic colors to simultaneously fit the multi-band light curves, the max_model approach avoids specifying host-galaxy extinction and instead focuses on extracting peak brightnesses across all observed filters (see also Stritzinger2010, their Sect. 3 for details).
The ‘max_model’ provides fit parameters of the time of maximum as ${\rm JD}\penalty 10000\ 2459316.74\pm 0.36$ and a color-stretch parameter $s_{BV}=1.19\pm 0.04$. Table 2 provides results from direct GP splines to the optical light curves, and the estimated peak apparent magnitudes of the optical/NIR light curves estimated from the SNooPymax_model’ fits, which include corrections for Milky Way reddening and computed K-correction. Adopting a redshift of $z=0.0363\pm 0.0049$ for SN 2021hem (see Sect. 3.4), the peak apparent magnitudes computed by the ‘max_model’ indicate that it reached an absolute $B$-band peak magnitude of $M_{B}=-19.90\pm 0.29$ mag, roughly half a magnitude brighter than typical SNe Ia of similar color-stretch values.
Table 2: Light curve fits.
Filter JD peak Peak app. Mag. $\Delta m_{15}$ (mag)
Gaussian Spline fits
$B$ $2459316.46\pm 0.31$ $16.338\pm 0.009$ $1.036\pm 0.051$
$g_{ztf}$ $2459316.60\pm 0.26$ $16.159\pm 0.014$ $0.823\pm 0.032$
$V$ $2459317.39\pm 0.37$ $16.250\pm 0.010$ $0.607\pm 0.034$
$r_{ztf}$ $2459316.07\pm 0.31$ $16.251\pm 0.016$ $0.469\pm 0.024$
$i_{ztf}$ $2459315.69\pm 0.32$ $16.698\pm 0.016$ $0.379\pm 0.028$
SNooPy ‘max_model fit’aaaaaaPeak magnitudes include K-correction and extinction correction. Uncertainties correspond to fit and systematic errors.
$B$ $\cdots$ $16.184\pm 0.011\pm 0.012$ $\cdots$
$g_{ztf}$ $\cdots$ $16.087\pm 0.013\pm 0.014$ $\cdots$
$V$ $\cdots$ $16.115\pm 0.015\pm 0.019$ $\cdots$
$r_{ztf}$ $\cdots$ $16.184\pm 0.012\pm 0.022$ $\cdots$
$i_{ztf}$ $\cdots$ $16.567\pm 0.026\pm 0.022$ $\cdots$
$J$ $\cdots$ $16.163\pm 0.077\pm 0.044$ $\cdots$
$H$ $\cdots$ $16.339\pm 0.062\pm 0.055$ $\cdots$
Inspection of the light curves also reveals several notable features suggesting SN 2021hem is a peculiar object. First, the flux evolution from the discovery epoch to the next observation shows a short plateau in the early light-curve and deviates markedly from a power-law rise, which is analyzed in detail in Sect. 4.1.2. Second, the $i$-, $J$-, and $H$-band light curves lack secondary maxima and appear to be unusually flat compared to what expected from a typical SNe Ia or luminous 1991T-like SNe Ia. This is clearly reflected in the significant deviations between SN 2021hem and the SNooPy template light curve fits in these bands, which are drawn from normal SNe Ia based on $\Delta m_{15}(B)$. Additionally, the NIR bands are nearly as bright as the optical, in contrast to normal SNe Ia, where NIR absolute magnitudes are typically about one magnitude fainter. Taken together, these photometric peculiarities of SN 2021hem, along with high peak luminosity, makes it resemble other 2003fg-like SNe Ia.
Figure 4: MCMC power-law model fit to ZTF $g$ and $r$-band fluxes to estimate the time of first-light $t_{first}$. The fitting is limited to the first 12 days after $t_{first}$, during which the light curve clearly follows a power-law rise. Power-law model fits are performed separately on the ZTF $g$- and $r$-band light curves to determine $t_{first}$ and $\alpha$ independently for each band. In the top panel, the $g$- and $r$-band fluxes are shown with arbitrary and distinct scales for visual clarity. The inferred (weighted mean) $t_{first}$ is shown by a vertical dashed line, with the shaded region indicating the uncertainty limits. In the bottom panels, the vertical and horizontal lines indicate the $t_{first}$ and $\alpha$ values, as inferred from each MCMC fit.
Ashall2020 proposed a photometric classification scheme for SNe Ia based on the time difference between the $i$- and $B$-band maxima ($t_{\max}^{i-B}$) and the color-stretch parameter $s_{BV}$. Using the $B$- and $i$-band peak dates listed in Table 2, we derive a rest-frame corrected value of $t_{\max}^{i-B}=-0.77\pm 0.45$ days for SN 2021hem. Combined with the previously estimated $s_{BV}$ range of 0.94–1.19, obtained from both direct $(B-V)_{0}$ color measurements and SNooPy fits, this places SN 2021hem among normal and luminous 1991T-like SNe Ia in the Ashall2020 classification diagram. In contrast, 2003fg-like subclasses occupy the region with $t_{\max}^{i-B}>0$ in the same diagram (Ashall2021). Therefore, SN 2021hem lies marginally outside the parameter space of 2003fg-like SNe sample that populate the Ashall2020; Ashall2021 classification diagram.
4.1.2 Estimating the time of first light, $t_{first}$
We fit the rising portions of the ZTF $g$- and $r$-band light curves individually using an MCMC power-law model (see Fig. 4), following the methodology of Bose2025. The time of first light, $t_{first}$, defined as the point when the SN luminosity begins a power-law rise, is estimated to be $16.63^{+0.71}_{-0.64}$ days relative to the epoch of $B$-band maximum, with a corresponding power-law index $\alpha$ of $1.52^{+0.14}_{-0.17}$ form the ZTF $r$-band light curve. Similarly, using the ZTF $g$-band light curve, the estimated parameters are $t_{first}=-16.31^{+0.58}_{-0.48}$ and $\alpha=1.54^{+0.10}_{-0.13}$. The parameter estimates from the individual light curves are highly consistent, and we adopt their weighted mean as $t_{first}=-16.43^{+0.45}_{-0.38}$ and $\alpha=1.53^{+0.08}_{-0.10}$. This time of first light corresponds to $\rm JD\penalty 10000\ 2459299.43^{+0.45}_{-0.38}$.
Interestingly, this epoch of $t_{first}$ occurs $1.51^{+0.45}_{-0.38}$ days after the initial detection (JD $\approx 2459297.92$) in the ZTF images. In fact, we re-ran the MCMC fit excluding the detection epoch, and found the epoch of first detection is outside 99.9% confidence level of $t_{first}$ estimation. Visual inspection of both the $r$- and $g$-band ZTF images confirms the presence of a source at this time, as shown in Fig. 3, marking the first detection in both the ZTF $g$ and $r$ bands at $\sim 20$ mags. This suggests that SN 2021hem emitted detectable radiation several days before its light curve began to follow the expected fireball expansion model.
The early SN atmosphere ($\lesssim 7$ day) can be well approximated as an optically thick “fireball” expanding homologously, and would result in a power-law flux rise ($\propto t^{\alpha}$). Assuming a constant temperature would imply a power-law index of 2. Any variation in the power-law index would imply a deviation of energy transport within the ejecta, primarily due to non-uniformity of 56Ni mixing. Studies implementing the fireball model with early SNe Ia light curves have shown significant deviations from the classical fireball model power-law index of 2. A study on a sample of early TESS SNe Ia light curves (2023ApJ...956..108F) found a mean power-law index of $1.93\pm 0.57$ with rise-time of $15.7\pm 3.5$ days, whereas another study using SNe Ia in PTF and LSQ survey data (2015MNRAS.446.3895F) found a mean power-law index of $2.44\pm 0.13$. In the case of SN 2020qxp/ASASSN-20jq, Bose2025 found a power-law index, which is almost exactly the same as the classical fireball model’s value of 2, indicating well-mixed 56Ni in the ejecta, which is also consistent with their non-LTE nebular-phase spectroscopic model that required macroscopic mixing to match observations.
Although the power-law index of $\penalty 10000\ 1.53$ for SN 2021hem lies within the range found in previous studies, it is on the lower side of the distribution and is significantly below the classical value of 2. The low power-law index of SN 2021hem possibly implies the SN had a shallow 56Ni mixing, where the 56Ni is confined deeper in the ejecta, thus the early light-curve rises more slowly because photons take longer to diffuse to the photosphere.
As discussed above, the first photometric detections of SN 2021hem were made several days prior to the onset of the light curve’s power-law rise. One explanation for this pre-fireball emission is shallow 56Ni mixing, where radioactive heating occurs deeper in the ejecta. This results in a ‘dark phase,’ during which $t_{first}$ is delayed by a few days after the explosion (see, e.g., 2013ApJ...769...67P; 2014ApJ...784...85P). This interpretation of the dark phase is consistent with the evidence for shallow 56Ni mixing discussed earlier in the context of the low power-law index. Another explanation for the delayed flux rise could be asymmetric distribution of 56Ni  where the majority of 56Ni is on the far side of the explosion. This would appear to have the same effect as deeply embedded 56Ni leading to a dark phase. An alternative interpretation, proposed for MUSSES1604D (2017Natur.550...80J), which displayed an early $\approx 1$ day plateau in its $g$-band light curve, involves a white dwarf undergoing a surface helium detonation, also referred to as an edge-lit detonation (e.g., 2010A&A...514A..53F; 2010ApJ...719.1067K). This surface detonation subsequently triggers a secondary detonation in the white dwarf core with a few days of delay, which can naturally explain the delayed fireball flux rise.
This pre-fireball emission should not be confused with the early flux excess seen in many early SNe Ia light curves, as the two phenomena are characteristically distinct. The early flux excess is typically attributed to the interaction of the SN ejecta with a non-degenerate companion star (2010ApJ...708.1025K), or with circumstellar material (e.g., Piro2016; Maeda2023; 2023MNRAS.522.6035M; 2018ApJ...864L..35S). It is estimated that approximately $20-30\%$ of all SNe Ia exhibit such early-excess features (2022MNRAS.513.3035M; 2022MNRAS.512.1317D). However, in certain subclasses of SNe Ia – such as the peculiar 2002es-like, 2003fg-like, and luminous 1991T-like events – the occurrence rate of early-excess emission appears to be considerably higher (2018ApJ...865..149J; 2024ApJ...966..139H). Crucially, unlike these early-excess signatures, the pre-fireball emission observed in SN 2021hem likely arises from a distinct physical mechanism, as discussed earlier. An early flux excess due to interaction typically emerges within a few days of explosion and appears superimposed on the power-law rise of the early light curve. This is because the short-lived excess flux from the interaction and the underlying “fireball” emission are both consequences of the expanding ejecta. In other words, the interaction can not occur without the “fireball” expansion. In contrast, the early emission seen in SN 2021hem occurs even before the onset of the fireball’s power-law flux rise, clearly distinguishing it from the early excesses or bumps arising from interaction in other SNe Ia. Moreover, 2017Natur.550...80J argued, based on their models, that an early excess arising from interaction would inevitably show a blue color evolution, as was observed in the $(\mbox{$B\!-\!V$})_{0}$ color of SN iPTF14atg. 2020ApJ...902...48B has also shown using ejecta-companion interaction models that the $(g-r)_{0}$ colors at the time of interaction is expected to be ${\sim}-0.5$ mag. In the case of SN 2021hem the $(g-r)_{0}$ color of $0.16\pm 0.33$ mag, computed from first epoch of $g$ and $r$-band light curves, does not show any blue excess. This color is broadly consistent with, or perhaps marginally redder than, the early $(g-r)_{0}$ colors typically observed in normal SNe Ia (2020ApJ...902...48B). Therefore, considering all these factors, the interaction scenario is disfavored for the pre-fireball emission seen in SN 2021hem.
Regardless of the origin, this early activity, which can be described as early flux excess or pre-fireball flux emission as in SN 2021hem, appears to be a ubiquitous feature among 2003fg-like SNe. In addition to SN 2021hem, other 2003fg-like supernovae with early-time observations and adequate photometric cadence – SNe 2020hvf, 2021zny, 2022ilv, and 2021qvo (2021ApJ...923L...8J; 2024ApJ...966..139H; 2025arXiv250813263A) – have all exhibited an early bump in their light curves.
Figure 5: UVOIR bolometric light curve of SN 2021hem fit with a radioactive 56Ni decay model. The time of explosion is taken as the epoch of first detection, which is $\approx 17.9$ days before the $B$-band maximum. Only data beyond 60 days are used for fitting the model. The UVOIR light curve of SN 2021hem (top panel), its time-weighted integrated luminosity (middle panel), and the ratio $t^{2}L/\left(\int Lt\penalty 10000\ dt\right)$ (bottom panel), which is defined independent of the 56Ni mass are shown.
4.1.3 Bolometric light curve and 56Ni mass
To estimate the quantity of 56Ni synthesized during the disruption of SN 2021hem, we constructed and modeled a UVOIR (ultraviolet–optical–infrared) bolometric light curve. Broad-band optical light curves were first interpolated using spline functions at coeval epochs. A total reddening correction of $E(B-V)_{\mathrm{tot}}=0.05$ mag was applied (see Sect. 3.3). The bolometric flux was then estimated using the SNooPy package, with spectral templates from 2007ApJ...663.1187H color-matched to the reddening-corrected broad-band colors of SN 2021hem. To account for flux outside the spectral range of the templates, we (i) extended the spectral energy distribution (SED) bluewards of the atmospheric cutoff by linearly extrapolating to zero flux at 1000 Å, and (ii) extended the redwards end of the SED beyond the $i$-band using a Rayleigh–Jeans tail extending to 25,000 Å. The integrated flux was then converted to bolometric luminosity using the distance adopted in Sect. 3.4. The resulting UVOIR bolometric light curve is shown in the top panel of Fig. 5.
To estimate the 56Ni mass, the UVOIR light curve of SN 2021hem was fit with an energy-deposition model corresponding to the radioactive decay chain $\mbox{${}^{56}$Ni}\rightarrow\mbox{${}^{56}$Co}\rightarrow\mbox{${}^{56}$Fe}$. The best-fitting model is overplotted on the UVOIR light curve in the top panel of Fig. 5. The model includes two free parameters: the synthesized 56Ni mass $M_{\mathrm{Ni}}$ and the $\gamma$-ray trapping timescale $t_{0\gamma}$. We assumed complete deposition of the positron kinetic energy from 56Co decay, while the fraction of trapped $\gamma$-ray energy was modeled as $\left[1-\exp\left(-t_{0\gamma}^{2}/t^{2}\right)\right]$. At late epochs, when the contribution from positron energy becomes comparable to that of the deposited $\gamma$-ray energy, degeneracy arises between $M_{\mathrm{Ni}}$ and $t_{0\gamma}$. To address this point, we examined the time-weighted integral of the luminosity (middle panel of Fig. 5). The result indicates that, at all times during the evolution of SN 2021hem, the energy budget can be fully explained by the 56Ni decay chain (Katz2013), with no indication from the model fit of any additional sources of energy input. The bottom panel of Fig. 5 shows the ratio $t^{2}L/\int Lt\,dt$, which is defined to be independent of $M_{\mathrm{Ni}}$, and therefore allows for an independent determination of $t_{0\gamma}$ (Katz2013). The model was fit only using data obtained beyond 60 days after explosion, when the ejecta is expected to be nearly optically thin and the luminosity is almost solely powered by radioactive decay. The best-fit parameters are $t_{0\gamma}=51\pm 3$ days and $M_{\mathrm{Ni}}=1.0\pm 0.1$ M. The inferred 56Ni mass is significantly larger than those typically found in normal SNe Ia (e.g., 2006A&A...450..241S; 2019MNRAS.483..628S; 2020MNRAS.496.4517S), but is consistent with larger values inferred for 2003fg-like SNe Ia (e.g., 2006Natur.443..308H; 2007ApJ...669L..17H; 2012ApJ...756..191K; 2019ApJ...880...35C).
Figure 6: Optical spectra of SN 2021hem from -10.2 days to ++33.8 days relative to the epoch of $B$-band maximum. Spectra are shifted by arbitrary constants for clarity. Telluric absorption features are marked by dashed lines and labeled with $\oplus$ symbols.
Figure 7: Comparison of similar phase spectra of the normal SN 2013gy and SN 2021hem. Telluric absorption features are marked by dashed lines and labeled with telluric symbols. Prominent spectral features in the pre-maximum light spectra are labeled.
4.2 Spectral properties, line diagnostics and comparative analysis
Fig. 6 presents our montage of seven optical spectra of SN 2021hem, spanning from 10.2-10.2 to +33.8+33.8 days. The spectra reveal characteristic SNe Ia features with the shape of the continuum evolving from the blue to the red over time as the time-series tracks the temperature evolution of the underlying emission regions. Close inspection of the 10.2-10.2 days spectrum reveals a prevalent C ii $\lambda$6580 feature appearing in the red wing of the most prominent feature Si ii $\lambda 6355$ (see line identifications in Fig. 7). Such a feature is estimated to occur in approximately 30% of SNe Ia (Parrent2011; Folatelli2012; Silverman2012; Maguire2014) and is typically associated with the presence of unburned carbon. However, it appears to be ubiquitous among 2003fg-like SNe Ia (e.g., 2007ApJ...669L..17H; 2011MNRAS.412.2735T; Ashall2021), which tend to have carbon-rich ejecta, and in some cases, it persists even beyond the time of maximum light (see below for more details). The measured pseudo-equivalent width ($pEW$) of the C ii feature from the 10.2-10.2 days spectrum is $\sim 8.9\pm 2.5$ Å. This is about a factor of 4 higher than that inferred from the normal SN 2011fe, and more consistent with values inferred from similar epoch spectra of the 2003fg-like SNe 2009dc and 2012dn (see Ashall2021; Lu2021). The measured velocity of C ii is $v_{abs}=13800\pm 560$$\rm{\,km\,s^{-1}}$ is also in agreement with measurements inferred from similar epoch spectra of the 2003fg-like objects including SN 2009dc, ASASSN-15hy, and ASASSN-15pz (Ashall2021; Lu2021). We also identify the feature near 7000  Å likely to be C ii $\lambda$7235 line, which is roughly at a velocity consistent with the C ii $\lambda$6580 line. Beyond the C ii features, the line features present in the time-series of SN 2021hem are consistent with many other normal SNe Ia.
Figure 7 shows a comparison of SN 2021hem spectra at representative phases with those of the spectroscopically normal SN 2013gy (Holmbo2019). The comparison includes SN 2021hem’s earliest spectrum at 10.2-10.2 days, one near maximum light (1.8-1.8 days), and another obtained approximately three weeks post-maximum. Key spectral features in the early-time spectra are labeled with their associated ions to aid in identification. Overall, SN 2021hem closely resembles SN 2013gy at all three epochs. The most notable difference is the presence of C ii absorption in the pre-maximum spectra of SN 2021hem, which is absent in SN 2013gy.
We now examine the standard spectroscopic diagnostics used for SNe Ia to determine where SN 2021hem fits within the established subtypes. This includes measuring the pEWs of the Si ii $\lambda$5972 and $\lambda$6355 features, which enables classification according to the Branch2006 scheme. We also measure the Doppler expansion velocity of the Si ii $\lambda 6355$ feature near maximum light, which provides insight into the ejecta kinematics.
The Branch diagram provides a quantitative framework for spectroscopic classification, relying on the pEWs of the Si ii $\lambda 5972$ and $\lambda 6355$ absorption features. Within this parameter space, SNe Ia are distributed into four principal subgroups: Core Normal (CN), Broad Line (BL), Cool (CL), and Shallow Silicon (SS), which reflect systematic variations in line strength and effective photospheric temperature (Nugent1995). The pEWs values of the Si ii $\lambda 5972$ and $\lambda 6355$ were measured from the 1.8-1.8 days spectrum by performing Gaussian fits, yielding values of $21.8\pm 1.8$ Å and $111.2\pm 1.0$ Å, respectively. These values place SN 2021hem firmly within the Branch diagram’s “core normal” subclass. 2003fg-like SNe are mostly associated with “shallow silicon” group in the branch diagram, while a few, for example, SNe 2012dn and CSS140501, are found to be associated with “core normal” group (Ashall2021). However, SN 2021hem has Si ii pEWs larger than any of these SNe, further highlighting the diversity among 2003fg-like SNe.
Using the SNIaDCA [^3] https://github.com/anthonyburrow/SNIaDCA code from Burrow2020, we find that SN 2021hem has an $85\%$ likelihood of belonging to the “core-normal” group. When additional components (i.e., $v_{\rm Si\penalty 10000\ II}$ and $M_{B}$) are included in the clustering analysis, the probability of SN 2021hem belonging to the “cool”-equivalent group increases dramatically. See Appendix C for further details on the probabilistic placement of SN 2021hem within these multi-dimensional clusters.
Figure 8: Si ii $\lambda$6355 velocity evolution of SN 2021hem compared with other 2003fg-like SNe Ia.
Following the Wang2009 framework, SNe Ia can be categorized as high-velocity (HV), with a blueshifted Doppler line velocities inferred from the position of maximum absorption (hereafter $v_{abs}$) of the Si ii line with $v_{abs}$ $\gtrsim 11{,}800$ km s-1, or otherwise “normal”, which populate a narrow strip in the Si ii velocity distribution, with an average $v_{abs}\sim 10{,}600\pm 400$ km s-1. Turning to SN 2021hem, we infer from the Si ii absorption component in the 10.2-10.2 days spectrum $v_{abs}\sim 12{,}500\pm 300$ km s-1. This decreases to $11{,}200\pm 400$ km s-1 by 1.8-1.8 days, and further declines to $9{,}200\pm 300$ km s-1 by +33.8 days. The velocity achieved around maximum is consistent with SN 2021hem being a normal SN Ia within the classification framework of Wang2009. Figure 8 compares the Si ii $\lambda$6355 velocity evolution of SN 2021hem with a sample of 2003fg-like SNe from Ashall2021. The velocity evolution is consistent with other 2003fg-like SNe, and particularly identical to SN 2012dn.
[…]
In summary, the spectroscopic line diagnostics of SN 2021hem are aligned with spectroscopically normal SNe Ia, except for the relatively prevalent C ii line present in our earliest optical spectrum, which is commonly seen in 2003fg-like SNe. The conclusion drawn from our near maximum-light spectra of SN 2021hem is consistent with many 2003fg-like SNe, as at this phase often these objects are spectroscopically indistinguishable from normal SNe Ia (see Sect. 6 of Ashall2021).
4.3 Revisiting redshift measurement and potential biases
It is important to note that all absorption-minimum velocities quoted in this section include an additional systematic uncertainty of $\Delta v_{\rm sys}=\pm 1{,}500\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$, propagated from the uncertainty in the redshift adopted in Sect. 3.4. 2003fg-like SNe exhibit a range of Si ii velocities, from relatively low values of $\sim 8{,}000\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$ (e.g. 2006Natur.443..308H; 2009ApJ...707L.118Y) to more typical SNe Ia-like values up to $\sim 12{,}000\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$ (e.g. 2007ApJ...669L..17H). Since the SN 2021hem redshift was estimated by matching SN Ia spectral templates using SNID (see Sect. 3.4), an intrinsically low-velocity SN could be matched to a normal-velocity template, thereby biasing the inferred redshift toward higher values. This, in turn, could lead to an overestimation of the absorption-minimum velocities from spectra corrected to rest-wavelength using the biased redshift. On the other hand, no significant bias is expected in template matching if the SN is intrinsically a normal velocity SN, which is on the higher end of 2003fg-like SNe velocities. Based on the analysis done so far, SN 2021hem appear to be an SN in the normal velocity range, yet, we are unsure if the measurements are biased by overestimated redshifts.
To investigate this potential bias, we examine the Si ii $\lambda 6355$ absorption line FWHM. The FWHM velocity ($v_{\rm FWHM}$) can be used as a tracer of the line velocity and is almost independent of the redshift correction applied to the spectrum. This may provide a clue whether SN 2021hem is intrinsically a normal-velocity or low-velocity 2003fg-like SN. Although the $v_{\rm FWHM}$ may not be mapped directly to absorption-minima velocity, it can serve as a proxy to relate with absorption-minima velocity. We measure both the Si ii $\lambda$6355 $v_{\rm abs}$ and $v_{\rm FWHM}$ for a number of 2003fg-like SNe spanning a range of velocities. As until the maximum light Si ii $\lambda$6355 line might be contaminated by the C ii line, we use coeval post-maximum spectra for this diagnostic. In the case of SN 2009dc at +4.5 days, $v_{\rm abs}=7{,}300\pm 150\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$ and $v_{\rm FWHM}=5{,}300\pm 200\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$; for ASASSN-15hy at +4.2 days, $v_{\rm abs}=7{,}900\pm 150\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$ and $v_{\rm FWHM}=5{,}900\pm 200\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$; for LSQ14fmg at +3.6 days, $v_{\rm abs}=10{,}400\pm 200\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$ and $v_{\rm FWHM}=6{,}300\pm 250\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$; and for SN 2012dn at +3.9 days, $v_{\rm abs}=10{,}500\pm 200\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$ and $v_{\rm FWHM}=6{,}400\pm 250\penalty 10000\ \mbox{$\rm{\,km\,s^{-1}}$}$. These values show a monotonic relation between $v_{\rm abs}$ and $v_{\rm FWHM}$. Now turning to SN 2021hem  the estimated velocities for +3.9 days spectrum are $v_{abs}=11,000\pm 350\mbox{$\rm{\,km\,s^{-1}}$}$ and $v_{\rm FWHM}=8{,}600\pm 250\mbox{$\rm{\,km\,s^{-1}}$}$. In the case of SN 2021hem, the measured $v_{\rm FWHM}$ is consistent with that of normal and higher-velocity 2003fg-like SNe. In fact, $v_{\rm FWHM}$ of SN 2021hem exceeds that of any of the measured 2003fg-like SNe, likely due to its extended blue-wing in the absorption line profiles, a feature not observed in any of the comparison objects. However, this analysis indicates that SN 2021hem is not an intrinsically low-velocity 2003fg-like SN, and therefore it is unlikely that the SNID-based redshift estimate is affected by a significant systematic overestimate.
5 Search for the host galaxy
One of the most intriguing aspects of SN 2021hem is its apparently hostless environment. SNe Ia are known to occur in a wide variety of galaxies and environments (2006ApJ...648..868S; Lampeitl2010; Pan2014), including stellar populations within galaxy clusters, which have been stripped from their hosts through tidal interactions and/or merger events. Only a handful of such intracluster SNe Ia have been identified to date (2003AJ....125.1087G; 2011ApJ...729..142S; 2015ApJ...807...83G). In the case of SN 2021hem, no nearby galaxy cluster is present in the field, making its origin from an intracluster star unlikely, although this possibility cannot be excluded. 2025ApJ...988..278S investigated the likely origins of SNe Ia that appear to lack associated host galaxies, particularly those occurring in open-field, non-clustered environments, and suggested that many of these seemingly hostless events are actually associated with faint dwarf galaxies that remain undetected in wide-field surveys due to shallow limited imaging depth.
This section examines whether SN 2021hem originated from a hyper-velocity progenitor that traveled far from its parent galaxy, or alternatively from an undetected low-luminosity host galaxy.
5.1 A runaway progenitor from its host
Unlike core-collapse SNe, SNe Ia exhibit long explosion delay times, typically ranging from $\sim 100$ Myr to several Gyr, with a delay-time distribution that follows approximately $t^{-1}$ (2008PASJ...60.1327T; 2012PASA...29..447M). In principle, such timescales allow a hyper-velocity progenitor to travel several hundred kiloparsecs from its host galaxy prior to explosion. To date, however, no confirmed case of a hyper-velocity progenitor has been identified.
Assuming SN 2021hem originated from a hyper-velocity runaway star (or binary system), we searched for known or visible galaxies in the vicinity of SN 2021hem as a plausible host. Two of the galaxies have been identified from SIMBAD (2000A&AS..143....9W) catalog as the closest and equidistant candidates from SN 2021hem (see Fig. 1). One is an elliptical galaxy LEDA 1461683 ($\alpha=16^{h}21^{m}25^{s}.051$; $\delta=+14^{\circ}34^{\prime}45\aas@@fstack{\prime\prime}60$) at $z=0.02873$, and a projected distance of $102\mbox{\penalty 10000\ kpc}$. The other is a LINER-type Active Galaxy Nucleus (AGN), 2MASX J16212572+1431537 ($\alpha=16^{h}21^{m}25^{s}.723$; $\delta=+14^{\circ}31^{\prime}53\aas@@fstack{\prime\prime}73$), which is at a similar projected distance of 104 kpc and redshift of $z=0.02952$. Although both of these galaxy redshifts differ from the plausible galaxy group’s redshift of $z\sim 0.035$ or the adopted SN redshift of 0.0363 as inferred in Sect. 3.4, such a discrepancy could be explained if the progenitor were a hyper-velocity star. A star with a line-of-sight velocity of $\sim 2000$ $\rm{\,km\,s^{-1}}$ would account for the higher observed redshift. As an interesting side note, considering the redshift and projected line-of-sight velocity of the star in this scenario, after accounting for reddening, luminosity distance, and K-correction, the absolute rest-frame magnitude of SN 2021hem would be $M_{B}=-19.45$ mag.
[…]
Figure 10: Image cutout centered on the position of SN 2021hem, taken on 9 September 2023 with the GTC equipped with OSIRIS. The final image is a stack of eight individual exposures, with a total integration time of 435 seconds. The location of the SN and the nearby diffuse extended feature is marked. An isophote is shown (green contour) enclosing the diffuse feature.
5.2 An undetectable faint dwarf host
We also explored the possibility that the SN resides in a low-luminosity dwarf host too dim to be detected in archival data. However, neither image analysis nor visual inspection revealed any trace of a host galaxy at the SN location, even in our deepest exposures. This makes it important to establish an upper limit on the host-galaxy brightness below which a source can not be detected in our data.
5.2.1 Determining the limiting magnitude from the GTC image
We use the 435s $r$-band GTC image shown in Fig. 10, obtained approximately 900 days after the explosion, to determine the limiting magnitude. After processing the image following the standard reduction procedure, image quality parameters: stellar FWHM and ellipticity are measured with SExtractor, and median sigma clipping is applied to reject nonstellar sources. Aperture photometry is then performed using the photutils DaoPhot package. The image is calibrated using field stars cross-matched with the PanSTARRS $r$-band catalog (Flewelling2020), and the zero-point magnitude for the image is determined to an uncertainty of $0.01$ mag.
[…]
Figure 11: Left: Detection efficiency as a function of magnitude with $3\sigma$ detection threshold. Each curve is based on 30,000 simulated images, generated by adding an artificial star drawn from a uniform distribution over the magnitude range and within a $\rm 4\cdot FWHM$ square region around the SN position. The different curves correspond to varying tolerances between the recovered flux and the true flux of the input star. A detection is counted as successful if the recovered flux lies within the specified tolerance level of the input artificial star. Right: Same as left panel, but using $1\sigma$ detection threshold and 50,000 simulations within magnitude range of 21.75 to 25.75 mag.
The detection efficiency curves are plotted in the left panel of Fig. 11, where each curve corresponds to different tolerances of recovered flux error with respect to the input artificial test star’s flux. For example, a 20% tolerance of recovered flux implies that a detection is considered true only if recovered flux is within 20% tolerance of its true flux of the artificial star. Similarly, the 99% tolerance curve implies the efficiency curve is almost solely determined if stars are detected above $3\sigma$ noise threshold. At the cutoff magnitude of $m_{cutoff,r}\approx 24.05$ mag, detection efficiency becomes zero for all the curves, as the test star is too faint to cross $3\sigma$ noise threshold at any location within $\sim 4\aas@@fstack{\prime\prime}7$ square area (as mentioned above).
Since we are primarily interested in detectability of a source (host galaxy) near the SN location, irrespective of the accuracy of the measured flux, we look into the 99% tolerance curve at 50% efficiency to determine the limiting-magnitude, and thus a value of $m_{lim,r}=23.8$ mag. This implies, assuming a point-source, any host associated with the SN is fainter than 23.8 mag, which translates to an absolute magnitude of 12.3-12.3 mag by adopting a luminosity distance for redshift of 0.0363.
It is important to emphasize that, in addition to image quality, the limiting magnitude we adopt also depends on the constraints imposed in our photometric workflow. The detection threshold is one of the primary parameters that governs the detectability of point sources. A second consideration is the definition of “sufficiently accurate” photometry, which we have already addressed through the generation of multiple efficiency curves for different tolerance levels (see Fig. 11), where each curve represents a different tolerance in accuracy of recovered flux. In most cases, a $3\sigma$ detection threshold provides a robust balance between accuracy and efficient detection, and this criterion was applied in the preceding analysis.
However, in this work, our primary interest lies in establishing the detectability of a possible host galaxy rather than obtaining reliable photometry of its features. For this purpose, a lower threshold is justified, and we find that a $1\sigma$ detection threshold represents a reasonable limit. For example, the faint diffuse feature highlighted in Fig. 10 (see Sect. 5.3) lies well below the $3\sigma$ noise level, but is detectable at $1\sigma$ level and also by visual inspection.
We therefore repeated the limiting-magnitude analysis using a $1\sigma$ detection threshold, with the results shown in the right panel of Fig. 11. The notable difference with a lower detection threshold is the fainter cutoff magnitude at $m_{cutoff,r}\approx 25.50$ mag, and the efficiency curves decline more gradually towards faint magnitudes. From the analysis with $1\sigma$ detection threshold, at 50% efficiency for the 99% tolerance curve, the limiting magnitude we find is $m_{lim,r}=24.4$ mag, or an absolute magnitude limit of 11.6-11.6 mag for any undetected galaxy.
5.2.2 Analytical estimate of limiting magnitude
As a sanity check for the limiting magnitude estimated above, we attempt to analytically calculate the limiting magnitude for the image based on the detection threshold. Assuming a Gaussian stellar PSF, the 2D profile is
$f(x,y)=h\cdot exp{\left[-\frac{(x-\bar{x})^{2}}{2\sigma_{x}^{2}}-\frac{(y-\bar{y})^{2}}{2\sigma_{y}^{2}}\right]},$ (1)
where $h$ is the amplitude of the Gaussian profile, and other symbols have their usual meanings. The volume of the 2D profile would be,
$V=h\cdot 2\pi\sigma_{x}\sigma_{y}.$ (2)
For a circular symmetric stellar PSF, $\sigma_{x}=\sigma_{y}=\sigma_{PSF}$. To satisfy a $3\sigma$ detection threshold, the amplitude $h$ must be $3\sigma_{sky}$, where $\sigma_{sky}$ is the local background sky-noise. Therefore, the limiting counts in ADU for the profile to satisfy $3\sigma$ detection threshold,
$C_{lim}=6\pi\sigma_{sky}\sigma_{PSF}^{2},$ (3)
and on converting to limiting magnitude,
$m_{lim}=ZP-2.5log_{10}\left[6\pi\sigma_{sky}\sigma_{PSF}^{2}\right],$ (4)
[…]
5.2.3 Limiting magnitude in terms of surface brightness
In the case of extended objects like galaxies, the detection limit can be better described in terms of surface brightness. Assuming a circular aperture area $A_{ap}$ within which the flux for the estimated point-source limiting magnitude is confined, the limiting surface-brightness can be approximated to be fainter by $\rm 2.5log_{10}(A_{ap})\,\mathrm{mag\penalty 10000\ arcsec^{-2}}$.
This limiting surface brightness should be regarded as an upper limit for the image. On considering a source that is extended over a larger area on the sky, the total integrated flux increases, while the noise per arcsec2 decreases by a factor of $1/\sqrt{Area}$ due to averaging. Consequently, for extended objects such as galaxies, and also depending on their morphology, it becomes possible to detect objects with fainter surface brightness. For surface brightness measured for a extended source over an area of $A_{sb}$, the S/N would scale up as $\sqrt{A_{sb}/A_{ap}}$, and the limiting surface-brightness would be fainter by a additional factor of $1.25\log_{10}(A_{sb}/A_{ap})\,\mathrm{mag\penalty 10000\ arcsec^{-2}}$. Therefore, the limiting surface-brightness would be:
$\mu_{lim}=m_{lim}+2.5\log_{10}(A_{ap})+1.25\log_{10}\left(\frac{A_{sb}}{A_{ap}}\right)\,\mathrm{mag\penalty 10000\ arcsec^{-2}},$ (5)
or
$\mu_{lim}$ $=m_{lim}+2.5\log_{10}\!\left(\tfrac{\pi}{4}\,{\rm FWHM}^{2}\right)$
$\quad+2.5\log_{10}\!\left(\frac{D_{sb}}{{\rm FWHM}}\right)\,\,\mathrm{mag\penalty 10000\ arcsec^{-2}},$ (6)
or
$\mu_{lim}=m_{lim}+2.5\log_{10}(\pi/4\penalty 10000\ {\rm FWHM}\cdot D_{sb})\,\mathrm{mag\penalty 10000\ arcsec^{-2}},$ (7)
where FWHM is the point-source FWHM in arcsec, and $D_{sb}$ is the diameter of the extended source in arcsec over which the surface-brightness is to be estimated. For our image, the equation reduces to $\mu_{lim}=m_{lim}+2.5\log_{10}(D_{sb})-0.094$.
[…]
There are a few caveats of this limiting surface-brightness estimate that should be noted. For a real galaxy, the limit of the surface brightness measurement would also be affected by the small-scale surface brightness variation of the galaxy. However, UDGs or UFDGs do not show much morphological diversity, and hence would not have a significant effect on our conversion. Secondly, when converting from a point source limiting magnitude to a surface-brightness limiting magnitude, we assume the sky background properties of the extended source, which contribute to detectability, are similar to those of the point source aperture and the sky annulus background. This also does not take into consideration the possible large variation in the sky background when an extended source area is being considered. However, this issue is somewhat mitigated, as in our point-source limiting magnitude simulation, we used a square area of $\sim 4\aas@@fstack{\prime\prime}7$ (3.8 kpc) width.
[…]
5.3 A nearby faint diffuse feature
[…]
If we assume the feature is a galaxy, the projected distance of SN at $\sim 5.5-6.4\mbox{\penalty 10000\ kpc}$ from the center of the feature (or $\sim 4.2\mbox{\penalty 10000\ kpc}$ from the edge) is unusually high for it to be the associated host, when the small size of this galaxy, or if a typical UGD, is considered. We look into the normalized directional light distance $d_{\mathrm{DLR}}$ of the SN (2006ApJ...648..868S; 2016AJ....152..154G), which is a measure of the distance between the SN and the host galaxy, normalized to the physical extent of the host galaxy (directional radius of host) in the direction of the SN. For SN 2021hem and this feature, we measure $d_{DLR}$ to be $3.1^{+1.2}_{-0.2}$. The uncertainty attributed to the measurement is primarily due to the irregular shape of the feature and the region of confusion arising in considering the center of the host galaxy for measurement. This measure essentially means, the distance to the SN is roughly three to four times farther than the visible radius of the host galaxy. We compared this with the $d_{\mathrm{DLR}}$ measurements of $\sim 1{,}600$ SNe from ZTF DR2 sample (2025A&A...694A..10D). The sample is cutoff to $z<0.06$, to ensure higher volumetric completeness, and also found the bias in $d_{\mathrm{DLR}}$ values towards higher redshift is minimal with this cutoff imposed as compared the full sample. The comparison reveals that only $\approx 1.8-4.0$ percentiles of SNe Ia in the sample have potential hosts that are further than the $d_{\mathrm{DLR}}$ value measured between SN 2021hem and this diffused feature. It is to be noted that in ZTF DR2 sample, the potential hosts are identified based on the closest ($d_{\mathrm{DLR}}$) galaxy candidate identified from Pan-STARRS1 and Legacy imaging surveys, and no attempt was made to detect hosts fainter than survey depths. Therefore, it is possible that the true percentile of hosts above the given $d_{\mathrm{DLR}}$ value could be even lower than inferred above.
Considering the large distance ($d_{\mathrm{DLR}}$) of the SN from the feature, and also the highly elongated and irregular shape of it, makes the diffuse feature a candidate of low probability to be the associated host of SN 2021hem. However, we cannot completely rule out that the diffuse feature is a galaxy, or if it is, its likelihood to be the associated host of SN 2021hem. If this is indeed the host of SN 2021hem, this would be the first SN detected in a UDG, and the progenitor has to be a high velocity star to reach this far before explosion. UGDs are low metallicity systems, and that favors being a host for 2003fg-like SNe. An alternative possibility could be that the diffuse feature is an unrelated background galaxy at a higher redshift.
6 Conclusions
The Type Ia SN 2021hem was discovered within 48 hours of last non-detection in ZTF survey, in an apparently hostless environment. Spectral template matching yields a redshift of $z=0.0363\pm 0.0049$, which is also consistent with an overdensity of similar redshift galaxies within $\sim$1 Mpc radius. Its peak absolute $B$-band magnitude of $M_{B,max}=-19.96\pm 0.29$ mag places SN 2021hem at the bright end of the SN Ia luminosity distribution. Unlike normal or luminous 1991T-like SNe Ia, however, the $i$- and NIR-band light curves show no secondary maximum – a hallmark of low-luminosity 1991bg-like SNe Ia and some other peculiar subtypes. The high peak luminosity combined with the lack of a secondary maximum in the $i$- and NIR-band light curves suggests SN 2021hem is a 2003fg-like SN Ia candidate.
The light curves of SN 2021hem exhibit slow evolution, characterized by shape parameters $\mbox{$\Delta m_{15}(B)$}=1.02\pm 0.02$ mag and $\mbox{$s_{BV}$}=0.94\pm 0.05$ mag. With these parameters, SN 2021hem is $\sim 0.3$ mag brighter than typical SNe Ia in both the luminosity–width and luminosity–color–stretch relations. By fitting the radioactive decay model to the tail of the bolometric light curve of SN 2021hem, we estimate a high 56Ni mass of $M_{Ni}=1.00\pm 0.09\penalty 10000\ M_{\odot}$, and a gamma-ray trapping timescale of $t_{0\gamma}=51\pm 3$ day. Such a 56Ni mass is significantly higher than that of normal SNe Ia, but consistent with a more massive progenitor as inferred for several other 2003fg-like SNe. The modeling of the time-weighted integral of the luminosity shows that the entire energy budget of the SN can be solely explained by the radioactive decay chain of 56Ni, without evidence of any significant additional energy source, like circumstellar interaction.
The Si ii velocity near maximum light is $\sim 11,200\mbox{$\rm{\,km\,s^{-1}}$}$, consistent with the normal-velocity class of SNe Ia as defined by Wang2009. The Si ii $\lambda$5972 and $\lambda$6355 line pEWs further place SN 2021hem within the core-normal subclass on the Branch diagram. Overall, the spectroscopic properties of SN 2021hem are consistent with those of normal SNe Ia. Notably, the earliest spectrum at 10.2-10.2 days reveals C ii $\lambda$6580 and $\lambda$7235 features at $\sim 13{,}800\mbox{$\rm{\,km\,s^{-1}}$}$. While the C ii absorption lines are not commonly observed in normal SNe Ia, it is particularly prevalent among 2003fg-like events. The slow-evolving light curves paired with luminosity, high inferred 56Ni mass, and the detection of C ii lines in early spectrum support the classification of SN 2021hem as a 2003fg-like SN Ia. Nevertheless, certain characteristics of SN 2021hem, such as its position within the normal SNe Ia parameter space ($t^{i-B}_{max}$ vs. $s_{BV}$) on the Ashall2020 classification diagram, and its relatively large Si ii pEWs compared to previously observed 2003fg-like SNe, indicate that SN 2021hem extends the observed diversity within this peculiar subclass.
On fitting a power-law model, as expected for a homologously expanding “fireball”, to the early $g$- and $r$-band fluxes yields a time of first light of $t_{\rm first}=-16.43^{+0.45}_{-0.38}$ days relative to $B$-band maximum, with a power-law index of $\alpha=1.53^{+0.08}_{-0.10}$. One particularly interesting aspect of SN 2021hem is that the first photometric detection occurs $1.51^{+0.45}_{-0.38}$ days before (with 99.9% confidence) the onset of the power-law flux rise. Unlike the flux excesses seen in other SNe Ia, this behavior cannot be explained solely by interaction with circumstellar material or a companion, which would inevitably produce an additional flux component superimposed on the fireball model’s power-law flux rise. Moreover, an emission arising from CSM interaction is expected to show a blue color, whereas in SN 2021hem the $(g-r)_{0}$ color of $0.16\pm 0.33$ mag does not show any blue excess, rather consistent with most typical SNe Ia early colors. A more plausible explanation for the observed pre-fireball emission is shallow 56Ni mixing, in which radioactive heating takes place deeper within the ejecta. This leads to a short “dark phase”, during which the fireball flux rise is delayed by a few days after the explosion. A shallow 56Ni mixing leading to slower radiation transport can also explain the smaller inferred value of power-law index $\alpha=1.53$, as compared to the classical “fireball” model’s value of 2. An asymmetric 56Ni distribution, which is concentrated on the far side of the explosion, may also lead to a similar dark phase.
In principle, a double-detonation model with an early helium flash can produce a short, 1122 day plateau in the early light curve, as observed in MUSSES1604D (2017Natur.550...80J). A thin shell of surface helium detonation may account for this early emission and subsequently trigger a detonation in the core of the WD with a few days of delay, thereby explaining the delayed onset of the fireball-like expansion light curve. However, a double-detonation scenario is deemed necessary only for sub-Chandrasekhar mass WDs. In the case of SN 2021hem, a 2003fg-like event, with a Chandrasekhar-mass WD already has a sufficiently high central density to initiate carbon ignition. Although it is theoretically possible for a double detonation with surface helium burning to occur in a Chandrasekhar-mass WD, no compelling observational evidence for such a process has been found to date. Instead, this would likely lead to an accretion-induced collapse into a neutron star rather than a thermonuclear explosion. We therefore regard this as a less likely scenario.
Regardless of the physical process behind this pre-fireball flux or early emission, all 2003fg-like SNe with sufficiently early observations – SNe 2020hvf, 2021zny, 2022ilv, and 2021qvo – exhibit similar early-time activity. SN 2021hem thus represents the fifth known event showing such behavior, underscoring the importance of prompt, high-cadence early observations to constrain the physical origin of early emission in these rare SNe Ia.
Another intriguing aspect of SN 2021hem is the absence of a clearly associated host galaxy. 2003fg-like SNe Ia discoveries tend to show a preference towards low-mass dwarf host galaxies (e.g., 2011ApJ...733....3C; 2013MNRAS.432.3117T), however, these SNe have been found in a range of host-galaxy types, from low-mass dwarfs to massive spiral galaxies (2025A&A...694A..10D; 2025MNRAS.542.2752B). To constrain the host and progenitor of SN 2021hem, various possibilities have been thoroughly investigated. The nearest identified galaxy is an AGN, and is located at a projected distance of 104 kpc, with a redshift of $z=0.02952$. Considering this as a plausible host, the progenitor could have been a hyper-velocity star (or binary system) ejected from it, traveling far off before the explosion. An AGN host can facilitate the production of a hyper-velocity progenitor. The transverse and radial velocity components of such a runaway hyper-velocity star can account for both the large offset from the host galaxy and the redshift difference between the host and that inferred for the SN. Assuming an explosion delay time of $100$ Myr for the SN, the estimated 3D unprojected velocity of this potential hyper-velocity progenitor is $\sim 2200$ km s-1.
The second possibility we explore is the presence of an undetected faint dwarf host. Deep GTC imaging obtained 2.5 years after the explosion revealed no host galaxy at the SN location. However, this observation allowed us to place an upper limit on any potential undetected host and to interpret it in the context of the galaxy population. At the $1\sigma$ detection threshold, the point-source limiting magnitude is estimated as $m_{lim,r}=24.4$ mag, and a limiting surface brightness of $\mu_{lim,r}=26.3$ mag arcsec-2. This surface brightness limit rules out the presence of nearly all galaxies within the known luminosity distribution, including most ultra-diffuse galaxies (UDGs), except for those at the faintest end of the UDG luminosity function ($26.3\lesssim\mu\lesssim 26.5$ mag arcsec-2). Nonetheless, the existence of an ultra-faint dwarf galaxy-like (UFDG) host cannot be excluded. Such galaxies, with their small sizes ($\sim 30$–70 pc) and extremely low surface brightness ($\mu>27$ mag arcsec-2), could remain undetected even beyond distances of just $\sim 20$ Mpc.
The deep GTC image reveals a faint, diffuse, and extended feature located at a projected distance of approximately $\sim 5.5-6.4$ kpc from the SN position. Its absolute magnitude is measured as $M_{r}=-13.2\pm 1.8$ mag, with a surface brightness of $\mu_{r}=25.58\pm 0.70$ mag arcsec-2. It remains uncertain whether this feature is a galaxy associated with the SN. Nevertheless, its measured surface brightness places it within the luminosity regime of UDGs. The feature, however, exhibits an unusually irregular and highly elongated morphology for a UDG in a cluster free environment. Furthermore, the normalized directional light distance ($d_{\mathrm{DLR}}$) of SN 2021hem from this feature is substantially larger than that of the vast majority of observed SNe. A comparison with the ZTF DR2 sample indicates that fewer than $4.0$ percentile of SNe may have potential hosts located at such large separations. Considering these aspects, if this feature is indeed a galaxy, it would be a candidate of significantly low probability, yet can not be ruled out, to be the associated host of SN 2021hem. On the other hand, UDGs are low-metallicity systems, which would be a favorable host for a 2003fg-like SNe. If this is the associated host, then SN 2021hem would be the first SN to be found in any UDG.
Taking into account all these factors, along with the limits we can place on any plausible undetected host galaxy, SN 2021hem remains one of the strongest candidates for a hostless supernova. Looking ahead, the combined capabilities of the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; 2019ApJ...873.. [^1] I), La Silla Quest Survey 4 (LS4; Miller2025), and The ATLAS Project (Tonry2018a) will greatly broaden the discovery space for such rare events.
Acknowledgements.
We are grateful to Subo Dong for suggesting our team to followup SN 2021hem, Mark Phillips for discussions, and Morgran Fraser for observational support. Funding of the Aarhus supernova group during the creation of this research comes from grants via the Independent Research Fund Denmark (IRFD; grant numbers 8021-00170B and 10.46540/2032-00022B), and the Aarhus University Research Foundation (Nova project AUFF-E-2023-9-28). L.G. and C.P.G. recognize the support from the Spanish Ministerio de Ciencia e Innovación (MCIN) and the Agencia Estatal de Investigación (AEI) 10.13039/501100011033 under the PID2023-151307NB-I00 SNNEXT project, from Centro Superior de Investigaciones Científicas (CSIC) under the PIE project 20215AT016 and the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, and from the Departament de Recerca i Universitats de la Generalitat de Catalunya through the 2021-SGR-01270 grant. C.P.G. acknowledges financial support from the Secretary of Universities and Research (Government of Catalonia) and by the Horizon 2020 Research and Innovation Programme of the European Union under the Marie Skłodowska-Curie and the Beatriu de Pinós 2021 BP 00168 programme. L.G. acknowledges financial support from AGAUR, CSIC, MCIN and AEI 10.13039/501100011033 under projects PID2023-151307NB-I00, PIE 20215AT016, ILINK23001, COOPB2304. Based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. The NOT data were obtained under programme ID P62-505 and P66-506. Based on observations collected at the European Organisation for Astronomical Research in the Southern Hemisphere, Chile, as part of ePESSTO+ (the advanced Public ESO Spectroscopic Survey for Transient Objects Survey – PI: Inserra). ePESSTO+ observations were obtained under ESO program ID 1103.D-0328. T.E.M.B. is funded by Horizon Europe ERC grant no. 101125877. J. D. acknowledges support by FCT for CENTRA through the Project No. UIDB/00099/2025. J. D. also acknowledges support by FCT under the PhD grant 2023.01333.BD.
Appendix A Broad-band Photometry of SN 2021hem
Optical and NIR broad-band photometry tables of SN 2021hem.
Table 3: Optical photometry of SN 2021hem.
UT Date JD - PhaseaaaaaaRest-frame days relative to the epoch of $B$-band maximum, i.e., JD 2459316.46. $B$ $V$ $g$ $r$ $i$ $c$ $o$ TelescopebbbbbbThe abbreviations for the telescope/instrument are: LCOGT - Las Cumbres Observatory telescope network observatons through FLOWS survey; NOT - Nordic Optical Telescope; ATLAS - ATLAS survey telescope’s orange and cyan filter observations; YSE - Pan-STARRS survey telescope observation through YSE survey; ZTF - ZTF survey telescopes. Data observed within 5 hr are represented under a single-epoch observation.
2,459,000 (days) (mag) (mag) (mag) (mag) (mag) (mag) (mag) / Inst.
2021-03-18.47 291.97 -23.6 ¿20.538 ¿20.140 ZTF
2021-03-22.39 295.89 -19.8 ¿19.600 ¿20.120 ZTF
2021-03-24.43 297.93 -17.9 20.132 $\pm$ 0.244 19.920 $\pm$ 0.211 ZTF
2021-03-27.50 301.00 -14.9 19.966 $\pm$ 0.263 19.656 $\pm$ 0.237 18.156 $\pm$ 0.780 ATLAS,ZTF
2021-03-28.42 301.92 -14.0 18.968 $\pm$ 0.139 18.913 $\pm$ 0.132 ZTF
2021-03-29.41 302.91 -13.1 18.543 $\pm$ 0.092 18.434 $\pm$ 0.076 ZTF
2021-03-30.46 303.96 -12.1 18.080 $\pm$ 0.058 18.036 $\pm$ 0.056 18.173 $\pm$ 0.079 ATLAS,ZTF
2021-03-31.43 304.93 -11.1 17.724 $\pm$ 0.051 17.846 $\pm$ 0.053 ZTF
2021-04-01.48 305.98 -10.1 17.401 $\pm$ 0.029 17.680 $\pm$ 0.048 17.436 $\pm$ 0.024 ATLAS,ZTF
2021-04-02.45 306.95 -9.2 17.059 $\pm$ 0.026 17.060 $\pm$ 0.032 ZTF
2021-04-03.43 307.93 -8.2 16.809 $\pm$ 0.025 16.924 $\pm$ 0.026 ZTF
2021-04-04.40 308.90 -7.3 16.733 $\pm$ 0.018 16.730 $\pm$ 0.029 17.137 $\pm$ 0.029 ZTF
2021-04-05.47 309.97 -6.3 16.548 $\pm$ 0.021 16.664 $\pm$ 0.021 16.776 $\pm$ 0.010 ATLAS,ZTF
2021-04-06.43 310.93 -5.3 16.427 $\pm$ 0.022 16.505 $\pm$ 0.027 ZTF
2021-04-06.49 310.99 -5.3 16.483 $\pm$ 0.008 YSE
2021-04-07.35 311.85 -4.4 16.761 $\pm$ 0.029 ZTF
2021-04-08.41 312.91 -3.4 16.284 $\pm$ 0.018 16.371 $\pm$ 0.024 ZTF
2021-04-09.32 313.82 -2.5 16.362 $\pm$ 0.016 16.321 $\pm$ 0.015 16.686 $\pm$ 0.020 LCOGT
2021-04-09.55 314.05 -2.3 16.293 $\pm$ 0.006 ATLAS
2021-04-10.06 314.56 -1.8 16.486 $\pm$ 0.033 16.251 $\pm$ 0.072 16.392 $\pm$ 0.012 16.759 $\pm$ 0.012 LCOGT
2021-04-10.44 314.94 -1.5 16.704 $\pm$ 0.028 ZTF
2021-04-11.36 315.86 -0.6 16.339 $\pm$ 0.019 16.239 $\pm$ 0.018 16.139 $\pm$ 0.029 16.042 $\pm$ 0.031 16.742 $\pm$ 0.028 LCOGT,ZTF
2021-04-12.45 316.95 0.5 16.326 $\pm$ 0.017 16.255 $\pm$ 0.017 16.766 $\pm$ 0.022 LCOGT
2021-04-13.41 317.91 1.4 16.210 $\pm$ 0.006 16.289 $\pm$ 0.006 16.719 $\pm$ 0.027 YSE,ZTF
2021-04-13.52 318.02 1.5 16.249 $\pm$ 0.006 ATLAS
2021-04-15.34 319.84 3.3 16.416 $\pm$ 0.020 16.274 $\pm$ 0.016 16.221 $\pm$ 0.018 16.369 $\pm$ 0.015 16.732 $\pm$ 0.024 LCOGT,ZTF
2021-04-15.49 319.99 3.4 16.425 $\pm$ 0.008 ATLAS
2021-04-17.44 321.94 5.3 16.743 $\pm$ 0.028 16.378 $\pm$ 0.007 ATLAS,ZTF
2021-04-17.58 322.08 5.4 16.331 $\pm$ 0.006 16.728 $\pm$ 0.008 YSE
2021-04-18.33 322.83 6.1 16.301 $\pm$ 0.023 16.343 $\pm$ 0.020 ZTF
2021-04-19.47 323.97 7.2 16.553 $\pm$ 0.009 ATLAS
2021-04-20.35 324.85 8.1 16.447 $\pm$ 0.021 16.404 $\pm$ 0.025 16.935 $\pm$ 0.028 ZTF
2021-04-22.46 326.96 10.1 16.575 $\pm$ 0.007 YSE
2021-04-23.11 327.61 10.8 16.887 $\pm$ 0.022 16.537 $\pm$ 0.016 16.756 $\pm$ 0.021 17.017 $\pm$ 0.027 LCOGT
2021-04-23.47 327.97 11.1 16.750 $\pm$ 0.014 ATLAS
2021-04-24.38 328.88 12.0 16.737 $\pm$ 0.023 16.636 $\pm$ 0.026 16.821 $\pm$ 0.020 ATLAS,ZTF
2021-04-25.01 329.51 12.6 16.568 $\pm$ 0.653 LCOGT
2021-04-25.41 329.91 13.0 16.743 $\pm$ 0.024 16.999 $\pm$ 0.035 16.848 $\pm$ 0.019 ATLAS,ZTF
2021-04-27.48 331.98 15.0 16.885 $\pm$ 0.061 ATLAS
2021-04-29.38 333.88 16.8 17.573 $\pm$ 0.062 16.949 $\pm$ 0.034 17.033 $\pm$ 0.038 16.906 $\pm$ 0.015 ATLAS,LCOGT
2021-04-30.42 334.92 17.8 17.245 $\pm$ 0.025 16.790 $\pm$ 0.026 17.074 $\pm$ 0.031 ZTF
2021-05-01.41 335.91 18.8 17.370 $\pm$ 0.031 16.869 $\pm$ 0.026 16.977 $\pm$ 0.031 ATLAS,ZTF
2021-05-02.40 336.90 19.7 17.404 $\pm$ 0.027 16.826 $\pm$ 0.026 ZTF
2021-05-02.61 337.11 19.9 16.907 $\pm$ 0.009 17.052 $\pm$ 0.010 YSE
2021-05-03.19 337.69 20.5 18.059 $\pm$ 0.029 17.175 $\pm$ 0.019 17.034 $\pm$ 0.024 LCOGT
2021-05-03.52 338.02 20.8 16.985 $\pm$ 0.012 ATLAS
2021-05-05.38 339.88 22.6 17.722 $\pm$ 0.030 16.896 $\pm$ 0.029 17.463 $\pm$ 0.065 ATLAS,ZTF
2021-05-06.40 340.90 23.6 17.886 $\pm$ 0.031 17.002 $\pm$ 0.024 ZTF
2021-05-07.36 341.86 24.5 17.970 $\pm$ 0.033 17.124 $\pm$ 0.022 ZTF
2021-05-07.58 342.08 24.7 17.107 $\pm$ 0.016 ATLAS
2021-05-08.37 342.87 25.5 18.066 $\pm$ 0.034 ZTF
2021-05-08.55 343.05 25.7 18.051 $\pm$ 0.015 17.198 $\pm$ 0.008 YSE
2021-05-09.46 343.96 26.5 18.122 $\pm$ 0.036 17.693 $\pm$ 0.022 ATLAS,ZTF
2021-05-10.36 344.86 27.4 17.151 $\pm$ 0.035 ZTF
2021-05-11.29 345.79 28.3 18.845 $\pm$ 0.106 17.707 $\pm$ 0.041 18.280 $\pm$ 0.037 17.232 $\pm$ 0.031 17.316 $\pm$ 0.076 LCOGT,ZTF
2021-05-11.49 345.99 28.5 18.297 $\pm$ 0.016 17.280 $\pm$ 0.008 17.318 $\pm$ 0.015 ATLAS,YSE
2021-05-12.38 346.88 29.4 18.936 $\pm$ 0.015 17.850 $\pm$ 0.012 18.367 $\pm$ 0.039 17.427 $\pm$ 0.010 LCOGT,ZTF
2021-05-13.47 347.97 30.4 17.876 $\pm$ 0.022 ATLAS
2021-05-14.29 348.79 31.2 18.922 $\pm$ 0.077 17.851 $\pm$ 0.064 17.469 $\pm$ 0.027 LCOGT
2021-05-15.37 349.87 32.2 18.513 $\pm$ 0.043 17.455 $\pm$ 0.031 17.422 $\pm$ 0.040 17.514 $\pm$ 0.018 ATLAS,ZTF
2021-05-17.41 351.91 34.2 18.088 $\pm$ 0.025 ATLAS
2021-05-18.34 352.84 35.1 18.597 $\pm$ 0.046 17.678 $\pm$ 0.033 17.530 $\pm$ 0.041 ZTF
2021-05-19.32 353.82 36.1 19.118 $\pm$ 0.042 18.121 $\pm$ 0.023 17.649 $\pm$ 0.032 17.767 $\pm$ 0.028 17.727 $\pm$ 0.014 ATLAS,LCOGT,ZTF
2021-05-19.51 354.01 36.2 18.606 $\pm$ 0.027 17.687 $\pm$ 0.014 YSE
2021-05-20.38 354.88 37.1 18.695 $\pm$ 0.053 17.750 $\pm$ 0.032 ZTF
2021-05-21.43 355.93 38.1 17.846 $\pm$ 0.039 17.827 $\pm$ 0.072 18.145 $\pm$ 0.030 ATLAS,ZTF
2021-05-22.34 356.84 39.0 18.596 $\pm$ 0.057 17.763 $\pm$ 0.036 ZTF
2021-05-23.40 357.90 40.0 17.883 $\pm$ 0.026 ATLAS
2021-05-24.37 358.87 40.9 18.667 $\pm$ 0.091 17.834 $\pm$ 0.046 17.835 $\pm$ 0.048 ATLAS,ZTF
2021-05-25.39 359.89 41.9 18.668 $\pm$ 0.107 17.854 $\pm$ 0.064 ZTF
2021-05-26.03 360.53 42.5 19.331 $\pm$ 1.136 17.952 $\pm$ 0.488 LCOGT
2021-05-27.31 361.81 43.8 18.803 $\pm$ 0.098 17.895 $\pm$ 0.046 ZTF
2021-05-27.45 361.95 43.9 18.788 $\pm$ 0.117 ZTF
2021-05-27.54 362.04 44.0 17.917 $\pm$ 0.168 ATLAS
2021-05-28.31 362.81 44.7 18.679 $\pm$ 0.076 17.893 $\pm$ 0.042 17.965 $\pm$ 0.056 ZTF
2021-05-29.35 363.85 45.7 18.755 $\pm$ 0.072 17.966 $\pm$ 0.040 ZTF
2021-05-29.40 363.90 45.8 18.137 $\pm$ 0.045 ATLAS
2021-05-29.50 364.00 45.9 18.030 $\pm$ 0.023 18.094 $\pm$ 0.023 YSE
2021-05-30.37 364.87 46.7 17.959 $\pm$ 0.041 18.115 $\pm$ 0.031 ATLAS,ZTF
2021-05-30.98 365.48 47.3 19.200 $\pm$ 0.086 18.379 $\pm$ 0.041 18.151 $\pm$ 0.038 LCOGT
2021-05-31.35 365.85 47.7 18.853 $\pm$ 0.059 18.020 $\pm$ 0.039 18.110 $\pm$ 0.056 ZTF
2021-06-01.34 366.84 48.6 18.922 $\pm$ 0.059 18.060 $\pm$ 0.041 ZTF
2021-06-02.36 367.86 49.6 18.874 $\pm$ 0.062 18.120 $\pm$ 0.042 18.209 $\pm$ 0.033 ATLAS,ZTF
2021-06-03.31 368.81 50.5 18.924 $\pm$ 0.058 18.161 $\pm$ 0.041 18.194 $\pm$ 0.059 ZTF
2021-06-04.31 369.81 51.5 19.471 $\pm$ 0.037 18.495 $\pm$ 0.025 18.895 $\pm$ 0.057 18.162 $\pm$ 0.042 18.391 $\pm$ 0.029 18.463 $\pm$ 0.026 ATLAS,LCOGT,ZTF
2021-06-06.34 371.84 53.4 18.987 $\pm$ 0.058 18.207 $\pm$ 0.042 ZTF
2021-06-06.41 371.91 53.5 18.275 $\pm$ 0.033 ATLAS
2021-06-06.50 372.00 53.6 18.942 $\pm$ 0.047 18.340 $\pm$ 0.022 YSE
2021-06-08.19 373.69 55.2 18.491 $\pm$ 0.076 ZTF
2021-06-09.47 374.97 56.5 19.102 $\pm$ 0.047 18.321 $\pm$ 0.025 YSE
2021-06-10.33 375.83 57.3 18.950 $\pm$ 0.060 18.398 $\pm$ 0.049 ZTF
2021-06-11.30 376.80 58.2 18.991 $\pm$ 0.062 18.364 $\pm$ 0.047 ZTF
2021-06-12.38 377.88 59.3 19.042 $\pm$ 0.062 18.417 $\pm$ 0.049 18.693 $\pm$ 0.033 ATLAS,ZTF
2021-06-13.19 378.69 60.1 19.543 $\pm$ 0.033 18.687 $\pm$ 0.025 18.696 $\pm$ 0.026 LCOGT
2021-06-13.28 378.78 60.1 19.044 $\pm$ 0.062 18.437 $\pm$ 0.050 18.582 $\pm$ 0.074 ZTF
2021-06-13.45 378.95 60.3 19.046 $\pm$ 0.030 18.584 $\pm$ 0.021 YSE
2021-06-14.33 379.83 61.2 19.051 $\pm$ 0.071 18.513 $\pm$ 0.052 ZTF
2021-06-15.08 380.58 61.9 19.569 $\pm$ 0.052 18.370 $\pm$ 0.208 LCOGT
2021-06-16.33 381.83 63.1 18.976 $\pm$ 0.105 18.580 $\pm$ 0.116 18.718 $\pm$ 0.113 ZTF
2021-06-16.45 381.95 63.2 19.105 $\pm$ 0.040 18.551 $\pm$ 0.031 YSE
2021-06-17.29 382.79 64.0 19.075 $\pm$ 0.087 18.511 $\pm$ 0.060 ZTF
2021-06-18.33 383.83 65.0 19.082 $\pm$ 0.097 18.543 $\pm$ 0.072 18.525 $\pm$ 0.048 ATLAS,ZTF
2021-06-18.81 384.31 65.5 19.614 $\pm$ 0.106 18.905 $\pm$ 0.054 18.873 $\pm$ 0.063 LCOGT
2021-06-19.43 384.93 66.1 19.211 $\pm$ 0.053 18.772 $\pm$ 0.026 18.868 $\pm$ 0.179 ATLAS,YSE
2021-06-20.37 385.87 67.0 19.138 $\pm$ 0.105 18.660 $\pm$ 0.071 18.739 $\pm$ 0.083 ATLAS,ZTF
2021-06-21.37 386.87 67.9 18.661 $\pm$ 0.109 ATLAS
2021-06-22.19 387.69 68.7 19.785 $\pm$ 0.236 18.804 $\pm$ 0.096 19.247 $\pm$ 0.163 18.593 $\pm$ 0.087 18.912 $\pm$ 0.132 LCOGT,ZTF
2021-06-22.43 387.93 69.0 18.967 $\pm$ 0.155 ATLAS
2021-06-23.38 388.88 69.9 18.900 $\pm$ 0.132 ATLAS
2021-06-24.36 389.86 70.8 18.942 $\pm$ 0.127 ATLAS
2021-06-24.87 390.37 71.3 19.573 $\pm$ 0.274 18.932 $\pm$ 0.136 18.785 $\pm$ 0.142 LCOGT
2021-06-25.40 390.90 71.8 18.409 $\pm$ 0.143 ATLAS
2021-06-26.31 391.81 72.7 18.803 $\pm$ 0.081 19.005 $\pm$ 0.134 ZTF
2021-06-26.38 391.88 72.8 18.886 $\pm$ 0.077 ATLAS
2021-06-26.41 391.91 72.8 18.811 $\pm$ 0.034 18.967 $\pm$ 0.032 YSE
2021-06-27.85 393.35 74.2 19.777 $\pm$ 0.127 19.109 $\pm$ 0.076 19.128 $\pm$ 0.087 LCOGT
2021-06-28.35 393.85 74.7 19.139 $\pm$ 0.080 18.927 $\pm$ 0.073 19.027 $\pm$ 0.061 ATLAS,ZTF
2021-06-29.25 394.75 75.5 18.964 $\pm$ 0.116 ZTF
2021-06-30.27 395.77 76.5 19.218 $\pm$ 0.159 18.739 $\pm$ 0.189 ZTF
2021-06-30.41 395.91 76.7 18.936 $\pm$ 0.031 19.139 $\pm$ 0.037 YSE
2021-07-02.35 397.85 78.5 19.406 $\pm$ 0.087 19.131 $\pm$ 0.134 19.106 $\pm$ 0.054 ATLAS,ZTF
2021-07-06.37 401.87 82.4 19.309 $\pm$ 0.034 19.327 $\pm$ 0.065 ATLAS,YSE
2021-07-08.38 403.88 84.4 19.227 $\pm$ 0.059 ATLAS
2021-07-10.34 405.84 86.2 19.237 $\pm$ 0.068 ATLAS
2021-07-13.40 408.90 89.2 19.516 $\pm$ 0.045 19.466 $\pm$ 0.042 YSE
2021-07-14.33 409.83 90.1 19.450 $\pm$ 0.070 ATLAS
Table 3: continued.
Table 4: NIR photometry of SN 2021hem.
UT Date JD - PhaseaaaaaaRest-frame days relative to the epoch of $B$-band maximum, i.e., JD 2459316.46. $J$ $H$ TelescopebbbbbbThe abbreviations for the telescope/instrument are: LCOGT - Las Cumbres Observatory telescope network observatons through FLOWS survey; NOT - Nordic Optical Telescope; ATLAS - ATLAS survey telescope’s orange and cyan filter observations; YSE - Pan-STARRS survey telescope observation through YSE survey; ZTF - ZTF survey telescopes. Data observed within 5 hr are represented under a single-epoch observation.
2,459,000 (days) (mag) (mag) / Inst.
2021-04-10.42 2459314.92 -1.5 16.409 $\pm$ 0.033 16.602 $\pm$ 0.044 NOT
2021-04-21.18 2459325.68 8.9 16.790 $\pm$ 0.027 16.628 $\pm$ 0.020 NOT
2021-05-11.22 2459345.72 28.2 16.953 $\pm$ 0.041 16.639 $\pm$ 0.032 NOT
Table 4: continued.
444
Appendix B
Figure 12: Intrinsic $(\mbox{$B\!-\!V$})_{0}$ color curve of SN 2021hem corrected only for Milky-Way reddening is shown, and is compared with normal SN Ia 2011fe, and low-luminosity SN Ia 1991bg.
Figure 13: Photometry region of the diffused extended feature at 4.2 kpc distance from SN location (yellow cross). The green polygon region is the aperture used for photometry, and the magenta polygon annulus represents the region for background estimation. The red contour represents the bright region of the feature used for estimating the surface brightness.
Appendix C Probabilistic Multi-dimensional Clustering Analysis
As previously discussed in Sect. 4, SN 2021hem possesses many of the characteristics of 2003fg-like SNe discussed in Ashall2021, (see their Fig. 18). As shown there, most of the 2003fg-like objects are classified as “shallow-silicon” objects within the Branch scheme (Branch2006), with a couple objects classified as “core-normals”. Burrow2020 re-formulate the Branch Diagram using Gaussian Mixture Models, providing a method by which to estimate the probability any individual SN Ia belongs to each Branch group. Their clustering analysis also extended beyond the Branch Diagram phase space (i.e., pEW Si ii $\lambda 5972$ vs. pEW Si ii $\lambda 6355)$ to include $v_{{\mathrm{Si\,\mathsc{ii}}}}$ and $M_{B}$ in both three- and four-dimensional spaces. While the Branch groups are robust and tightly constrained even in higher-dimensional spaces, the groupings in these higher-dimensional spaces are by definition distinct from the original Branch groups.
Using the measurements from Ashall2021 and those of SN 2021hem determined in Sect. 4.2, we re-plot the traditional (e.g. two-dimensional) Branch Diagram in Fig. 14. Because group membership is determined probabilistically, we find that all 2003fg-likes now belong to the “shallow-silicon” group, with a minimum membership probability of $p>0.79$ (see Table 5 for full probabilities). We find SN 2021hem is classified as a “core-normal” with $p=0.8526$, significantly offset from the rest of the 2003fg-like sample. This is consistent with the placement of SN 2021hem in the $s_{BV}$ vs. $t^{i-B}_{max}$ space defined in Ashall2021 where it sits outside the region of 2003fg-like SNe Ia.
Figure 14: Branch Diagram including SN 2021hem (assuming $z=0.0363$), the sample of 2003fg-like SNe Ia from Ashall2021 over-plotted with the Burrow2020 sample. We note that both CSS140501 and SN 2012dn have both moved from the “core-normal” classifications assigned in Ashall2021 to the “shallow-silicon” groups here. See Table 5 for the full probabilities, which correspond to the colorings of the overplotted points.
Table 5: Branch Group Membership Probabilities of 2003fg-like SNe Ia
SN Name CN $S$ BL CL
SN 2003fg ${4.147}\times 10^{-4}$ 0.9996 ${1.207}\times 10^{-16}$ ${3.758}\times 10^{-6}$
SN 2006gz ${6.572}\times 10^{-4}$ 0.9993 ${5.780}\times 10^{-17}$ ${3.908}\times 10^{-6}$
SN 2012dn 0.2025 0.7962 ${2.515}\times 10^{-12}$ 0.0013
SN 2007if ${2.490}\times 10^{-9}$ 0.9999 ${2.834}\times 10^{-34}$ ${1.453}\times 10^{-10}$
SN 2009dc ${9.233}\times 10^{-5}$ 0.9999 ${6.830}\times 10^{-21}$ ${5.998}\times 10^{-7}$
LSQ12gpw 0.0048 0.9952 ${2.368}\times 10^{-14}$ ${2.594}\times 10^{-5}$
SN 2013ao ${1.376}\times 10^{-4}$ 0.9999 ${1.906}\times 10^{-21}$ ${8.968}\times 10^{-7}$
CSS140501 0.1244 0.8748 ${3.976}\times 10^{-16}$ ${8.772}\times 10^{-4}$
LSQ14fmg ${8.573}\times 10^{-7}$ 0.9999 ${2.002}\times 10^{-27}$ ${1.273}\times 10^{-8}$
SN 2015M ${1.272}\times 10^{-5}$ 0.9999 ${6.699}\times 10^{-21}$ ${1.606}\times 10^{-7}$
ASASSN15hy ${4.136}\times 10^{-7}$ 0.9999 ${2.955}\times 10^{-30}$ ${9.391}\times 10^{-9}$
SN 2021hem 0.8526 ${3.751}\times 10^{-4}$ ${5.936}\times 10^{-6}$ 0.1470
Extending the above analysis to higher dimensional spaces by including $v_{{\mathrm{Si\,\mathsc{ii}}}}$ and $M_{B}$ in the GMM clustering requires addressing the uncertainties in both quantities associated with the seeming hostless nature of SN 2021hem. We examine 3 cases: (1) assuming SN 2021hem lies at $z=0.0363$ as adopted in the main body of the paper; (2) the runaway progenitor case which assuming SN 2021hem lies within 2MASX J16212572+1431537 at $z=0.02952$; and (3) assuming $z=0.0363$ but with larger uncertainties derived from propagating the redshift uncertainty as discussed in Sect. 3.4. We note that cases (1) and (2) consider only the measurement uncertainties associated with the Si ii pEWs, $v_{\rm Si\penalty 10000\ II}$, and $M_{B}$. The results of this analysis are shown below in Fig. 15 and Table 6.
We find that that despite the overluminous nature of SN 2021hem, the inclusion of $M_{B}$ in both cases (1) and (3) changes the membership of SN 2021hem from the “core-normal”-equivalent group to the “cool”-equivalent group. The inclusion of $v_{\rm Si\penalty 10000\ II}$ causes the probabilities to shift 20% in the same direction without changing the group membership of SN 2021hem. Given the well-established association between underluminous, rapidly-declining SNe Ia and the “Cool” Branch group (Branch2006; Burrow2020; Phillips2025); the placement of the overluminous, slowly-declining SN 2021hem in the “cool” group is likely a reflection of the abnormal, hostless nature of SN 2021hem, but may hint at our incomplete understanding of of 2003fg-like SNe Ia. This placement likely arises due to: (1) a tight clustering of the “core-normal” objects along the $v_{\rm Si\penalty 10000\ II}$ and $M_{B}$ axes, (2) an upper-limit on pseudo-equivalent widths of the Si II lines for membership in the “shallow-silicon” group, and (3) the “broad-line” group being defined by high photospheric velocities to the point of being almost completely distinct from the other Branch groups in the higher dimensional spaces (Burrow2020).
In case (2), we find that SN 2021hem is associated with the “core-normal”-equivalent group at $p>0.8$ in all dimensionalities. SN 2021hem’s placement outside of the “shallow-silicon” (or equivalent) group in all cases and dimensionalities suggests that within this clustering scheme, its properties are significantly different than expected for a 2003fg-like SNe Ia at $z=0.0363$. Instead, this method of analysis, taken in combination with the location of SN 2021hem in the Ashall2020 classification, is more consistent with SN 2021hem being a overluminous “core-normal” SN Ia arising from the hyper-velocity progenitor scenario discussed in Sect. 5.1. However, given both the diversity of 2003fg-like SNe Ia behavior and the hostless nature of SN 2021hem, such results should be evaluated in combination with other analyses and remembering the large uncertainties on all measurements.
Figure 15: Multi-dimensional clusterings projected into the Branch Diagram phase-space for all combinations of input data and cluster dimensionality. The full probabilities associated with the group membership of SN 2021hem can be found in Table 6. Numerical labels on the overplotted data are the same as in Fig. 14.
Table 6: Multi-dimensional “Branch” Group Membership Probabilities of SN 2021hem
CN $S$ BL CL
Case 1: $z=0.0363$
2D (Branch Diagram) 0.8526 0.0004 ${5.936}\times 10^{-5}$ 0.1470
[…]
Case 2: Runaway Progenitor ($z=0.02952$)
2D (Branch Diagram) 0.8018 ${8.450}\times 10^{-5}$ ${3.123}\times 10^{-6}$ 0.1981
[…]
Case 3: $z=0.0363$ with propagated uncertainties
2D (Branch Diagram) 0.8295 0.001 ${4.318}\times 10^{-6}$ 0.1704
[…]